### v1

# Aerial Vision-and-Language Navigation via Semantic-Topo-Metric Representation Guided LLM Reasoning

Abstract
Aerial Vision-and-Language Navigation (VLN) is a novel task enabling Unmanned Aerial Vehicles (UAVs) to navigate in outdoor environments through natural language instructions and visual cues. It remains challenging due to the complex spatial relationships in outdoor aerial scenes. In this paper, we propose an end-to-end zero-shot framework for aerial VLN tasks, where the large language model (LLM) is introduced as our agent for action prediction. Specifically, we develop a novel Semantic-Topo-Metric Representation (STMR) to enhance the spatial reasoning ability of LLMs. This is achieved by extracting and projecting instruction-related semantic masks of landmarks into a top-down map that contains the location information of surrounding landmarks. Further, this map is transformed into a matrix representation with distance metrics as the text prompt to the LLM, for action prediction according to the instruction. Experiments conducted in real and simulation environments have successfully proved the effectiveness and robustness of our method, achieving 15.9% and 12.5% improvements (absolute) in Oracle Success Rate (OSR) on AerialVLN-S dataset.
I Introduction
The Aerial Vision-and-Language Navigation (Aerial VLN) [liu_2023_AerialVLN] emerges as a groundbreaking task. It enables unmanned aerial vehicles (UAVs) to interpret natural language instructions and corresponding visual information to navigate in outdoor environments. This technology can eliminate the necessity for manual UAV operation by human pilots, clearly mitigating the barriers to human-UAV interaction and potentially benefitting search, rescue, and delivery tasks.
Recently, VLN tasks have been well-developed. Considering the powerful reasoning capability of large language models (LLMs) [ahn2022can, song2023llm, driess2023palm, huang2022inner], several VLN methods have started to use LLMs as agents to parse instructions and predict actions [zhou2024navgpt, shah2023lm, li2024tina, lin2024navcot]. Specifically, existing methods attempt to describe the visual observations with text to enhance the LLM’s scene understanding ability. NavGPT [zhou2024navgpt] converted visual scene semantics into text description for the LLM to conduct high-level planning. MapGPT [chen2024mapgpt] introduced a topological graph described in language as text prompts to an LLM, guiding low-level planners to explore new nodes of a predefined graph of the environment.
Figure 1: The pipeline to obtain STMR. Observation image, semantic mask, and depth image (a) will be aggregated in the top-down map (red dashed box) (b), which captures the Semantic-Topo-Metric information of the neighborhood of the current position, further generating a 20x20 matrix representation (c).
Although existing VLN methods have made significant progress, most of them are designed for indoor or ground-based outdoor environments and pay less attention to aerial VLN tasks. Practically, the rich semantics, complex spatial relationships, and the large-scale nature of outdoor 3D spaces make it challenging for existing VLN methods to adapt to aerial scenarios. As shown in  Fig. 1 (a), the aerial scene can be quite complex, which may lead to incorrect reasoning for several reasons, such as an overemphasis on instruction-irrelevant objects, or a failure to capture the contextual relationships between different areas. Furthermore, the vast scale of aerial landscapes can further complicate navigation, making local-perspective-based VLN methods unable to comprehend the environment. Thus, the development of VLN methods that integrate both effective semantic information and precise spatial representations is urgently needed to enhance the adaptability and accuracy in aerial scenarios.
To overcome these challenges, we propose a zero-shot LLM-based aerial VLN framework, which takes natural language instructions, RGB images, and depth images as input, and directly makes action predictions (e.g., forward 10 meters) through single-step reasoning updates. Specifically, a Semantic-Topo-Metric Representation (STMR) is designed for LLMs prompting. We first extract instruction-related landmarks and obtain corresponding semantic masks by perception models, i.e., Grounding Dino [liu2023grounding] and Tokenize Anything [pan2023tap]. After that, semantic masks obtained from each step are gradually projected into a top-down map using depth and pose transformation as shown in Fig. 1 (b). This projected map encompasses both the UAV’s travel trajectory and spatial information, i.e., locations of surrounding objects. To integrate rich visual information and topology into the text prompt while maintaining its simplicity, we separate the top-down map centered at the UAV’s current position into grids and substitute each grid with a semantic number. This process transforms the top-down map into a matrix representation, tailored to serve as a spatial prompt for the LLM. As can be seen from Fig. 1 (c), the matrix representation encompasses topological, semantic, and metric information, and is input to an LLM together with historical actions, sub-goal instructions, and other information to infer the next action. Compared with other learning-based and LLM-based methods, the proposed method effectively improves the accuracy of spatial reasoning.
Our contributions are summarized as follows:
  • To our knowledge, we propose the first LLM-based end-to-end framework for the aerial VLN task, facilitating the development of UAV navigators. This framework requires no training and no extra low-level action planner, thus allowing for easy integration.
  • We propose the Semantic-Topo-Metric Representation (STMR), a unique matrix representation that encompasses topological, semantic, and metric information. STMR is designed to significantly enhance the spatial-aware reasoning capabilities of LLMs in outdoor environments.
  • Extensive experiments on the aerial VLN task especially in real environments demonstrate that the proposed method outperforms previous state-of-the-art methods by a large margin, establishing a strong baseline for future zero-shot Aerial VLN tasks.
II Related Work
Vision-Language Navigation (VLN)
VLN aims to enable autonomous agents to navigate complex environments by understanding and executing natural language instructions based on visual context. Early VLN methods use sequence-to-sequence LSTMs to predict low-level actions [anderson2018vision] or high-level actions from panoramas [fried2018speaker]. Several attention processes have been proposed [qi2020object, hong2020language, an2021neighbor] to enhance the process of learning visual textual correspondence. Reinforcement learning is also explored to improve policy learning [wang2018look, tan2019learning, wang2020soft]. Besides, transformer-based architecture have shown superior performance to long-distance contextual information [hao2020towards, majumdar2020improving]. More recent works [zhou2024navgpt, chen2024mapgpt] leverage the reasoning and dialogue capabilities of LLMs, achieving great progress. However, most of them are developed under a ground-based, discrete VLN setting, which simplifies navigation as traversing on a predefined graph on the ground. This limits the free movement space of UAVs in the real world. Aerial VLN remains challenging due to the large-scale and complex environments.
UAV navigation
Unmanned Aerial Vehicle (UAV) navigation has seen a surge of interest over the past few decades. Guided with instructions, Blukis et al. [blukis2018following] use imitation learning for low-level velocity command prediction. [blukis2018mapping] predicts interpretable position-visitation distributions to guide UAV’s control actions, using a combination of supervised and imitation learning for efficient training. Misra et al. [misra2018mapping] decompose instruction execution in two stages, separately using supervised learning for goal prediction and policy gradient for action generation. AerialVLN [liu_2023_AerialVLN] contributes a much more challenging aerial VLN dataset focusing on large-scale environments, and provides a look-ahead guidance method as the baseline. Despite the progress, the generalizability and performance of these methods still require improvement.
LLMs for robot planning and interaction
Most recently, LLMs have demonstrated impressive capabilities in understanding and reasoning. To leverage these capabilities, several promising methods have been proposed for applying LLMs in robotic systems. A few methods involve using LLM-generated rewards optimized in simulation to improve control [huang2023voxposer, yu2023language]. Others utilize LLM-selected subgoals as an abstraction to enhance policies for navigation [dorbala2022clip, chen2023open] and manipulation [cui2022can, shridhar2022cliport]. Additionally, research has explored the use of LLMs to generate executable code for control and perception primitives [singh2023progprompt, wu2023tidybot, liang2023code]. Despite their potential, LLMs are still prone to confidently hallucinating outputs, such as referring to objects not observed in the scene [zeng2022socratic]. In order to alleviate the hallucinating phenomenon, we propose a matrix-based representation containing topological, semantic, and metric information for better prompting the LLMs.
III Method
In this paper, we introduce a novel zero-shot framework that leverages large language models (LLMs) for action prediction in aerial VLN tasks. As shown in Fig. 2, our framework mainly consists of three modules. The sub-goal extracting module decomposes language instructions into several sub-goals, facilitating step-by-step reasoning and navigation. The Semantic-Topo-Metric Representation (STMR) module represents the outdoor environment as a matrix containing semantic, topological, and metric information for the prompt to an LLM. Finally, the LLM planner outputs its current thoughts and actions by using the prompts from the aforementioned two modules along with the task description and history as inputs.
III-A Problem Formulation
In an aerial VLN task, the problem is formulated as a free-form language instruction guided navigation. At the beginning of each episode, the UAV is placed in an initial pose $P=[x,y,z,\phi,\theta,\psi]$, where $(x,y,z)$ denotes the UAV’s position and $(\phi,\theta,\psi)$ represents pitch, roll, and yaw of the UAV’s orientation. A natural language instruction $L$ is provided to specify the path that a UAV should follow. To achieve the navigation goal, the UAV considers both the instruction and visual perceptions, and predicts an action from the action space (right, left, up, down, forward, backward) with corresponding parameters at each time step $t$. Navigation ends when the UAV predicts a ’Stop’ action or reaches a pre-defined maximum action number. The navigation is recognized as a success if the UAV stops at a location that is less than 20 meters from the target location.
Figure 2: Our method consists of three modules, i.e. sub-goal extracting, Semantic-Topo-Metric Representating, and LLM planner. First, we prompt an LLM to break down long instructions and extract specific landmarks, which will be classified by perception models, generating corresponding semantic masks. Then we project these masks into a top-down map using the UAV-view pose and depth information. The map is further compressed into a matrix representation. Finally, we prompt LLM with well-designed instructions to enable effective reasoning.
III-B Semantic-Topo-Metric Representation (STMR)
Previous LLM-based VLN methods use natural language to describe current observations, or nodes and edges of a topological graph to model the spatial information of the environment. However, in open scenarios, simple directional words such as "next to" or "aside" are not enough to describe complex spatial relationships, which can easily introduce ambiguity into LLMs. To address this challenge, we introduce the Semantic-Topo-Metric Representation (STMR) to enhance the spatial-aware reasoning capability of LLMs. Specifically, STMR incrementally takes an RGB image $I_{t}^{R}$ and a depth map $I_{t}^{D}$ as input from each step, and generates a dynamically updated matrix representation with semantic, topological, and metric information as its output. The details of STMR are presented as follows.
2D Visual Perception
Impressed by the powerful open-vocabulary detection capabilities of Grounding DINO $\operatorname{GD}(\cdot)$, as well as the captioning and segmentation capabilities of Tokenize Anything model $\operatorname{TA}(\cdot)$, we integrate these two models as our 2D visual perceptor, as illustrated in Fig. 3. Given a single RGB image $I_{t}^{R}$ and an instruction $L$ as input, we first obtain detailed landmark categories using a Landmark Extractor $\operatorname{LE}(\cdot)$ driven by an LLM, and then identify the corresponding bounding box for each category through $\operatorname{GD}(\cdot)$. Next, we employ $\operatorname{TA}(\cdot)$ to take each bounding box as a prompt and output a set of 2D semantic masks $m^{(t)}$ and captions $z^{(t)}$ for the current RGB image $I_{t}^{R}$. The entire process can be described as:
Figure 3: 2D Visual Perceptor for the UAV.
$\{m^{(t)},\ z^{(t)}\}=\operatorname{TA}\left({I_{t}^{R}},\ \operatorname{GD}\left({I_{t}^{R}},\ \operatorname{LE}({L})\right)\right).$ (1)
To improve the simplicity of the semantic masks and reduce the misleading of LLM-reasoning caused by numerous open-vocabulary categories, we propose a text-matching method to mitigate over-classification. As shown in Fig. 3, our method involves vectorizing the landmarks extracted from the instruction and the captions $z^{(t)}$ generated in each $I_{t}^{R}$. Then, we calculate the cosine similarity between these vectorized landmarks and captions using TF-IDF [qaiser2018text]. If the similarity score exceeds the threshold $\tau$ >0.8, the landmark is classified as visible in the current view. By implementing this strategy, we effectively simplify semantic masks and ensure the LLM’s reasoning focuses on relevant categories.
Sub-goal-driven top-down map
Considering that the top-down view can better express spatial relationships than the first-person view, we further utilize the depth image $I_{t}^{D}$ to map semantic masks to 3D space, and then convert them into a top-down map. Specifically, with the RGB image $I_{t}^{R}$ processed by the 2D visual perceptor, we produce first-person view predictions by assigning semantic labels to each pixel, resulting in a segmented image with identified objects and regions. Subsequently, the depth image $I_{t}^{D}$ is converted into a 3D point cloud, where each pixel is mapped to a 3D point $(X,Y,Z)$ based on its depth value and camera parameters:
$X=\frac{(u-c_{x})I_{t}^{D}(u,v)}{f_{x}},Y=\frac{(v-c_{y})I_{t}^{D}(u,v)}{f_{y}},Z=I_{t}^{D}(u,v),$ (2)
where $(u,v)$ are the pixel coordinates, $(c_{x},c_{y})$ are the camera’s principal point coordinates, and $(f_{x},f_{y})$ are the focal lengths. The semantic labels from the segmented RGB image are mapped to the corresponding 3D point clouds, resulting in a point cloud with semantic information $(X,Y,Z,C_{i})$, where $C_{i}$ are the semantic categories. Then the 3D point cloud is partitioned into discrete voxels, where each voxel aggregates its point clouds as one semantic category using max pooling described in [chaplot2020learning]. Considering that UAVs usually fly above ground objects, for a specific coordinate $(x,y)$, the semantic label of the top one in a column of voxels will be projected into the top-down map:
$\text{TopDownMap}(x,y)=\text{Voxel}(x,y,z_{top}),$ (3)
where $z_{top}$ means the highest $z$ coordinate at location $(x,y)$ and $\text{Voxel}(x,y,z_{top})$ denotes the corresponding semantic label. So, we get a top-down map with semantic information. Notably, since the LLM prioritizes the landmarks in the current sub-goal, if a category contained in the sub-goal appears in the voxel at any $z$ coordinate, this category will be projected into the top-down map first. Thus we can modify Eq. 3 as:
$\text{TopDownMap}(x,y)=\begin{cases}C_{i},&C_{i}\text{ in sub-goal}\\ \text{Voxel}(x,y,z_{top}),&\text{otherwise}.\end{cases}$ (4)
At each step, the position of a UAV will also be recorded as waypoints with special values and projected onto the top-down map.
Matrix representation
We found that directly inputting the image-format top-down map to a vision-language model (VLM) often yields poor reasoning results. Therefore, we process the visual map into a text-based matrix representation, which is input to an LLM to guide the navigation. Specifically, the UAV incrementally collects observations and gradually constructs the sub-goal-driven top-down map of the surrounding environment. Considering the perception quality of onboard sensors and the effective processing capacity of modern LLMs, we define a local map centered at the UAV’s current position as shown in Fig. 1. Subsequently, we generate a matrix by partitioning the local map into a $20\times 20$ coordinate grid and applying semantic max pooling in each grid, making them finally contain one semantic category. Particularly, we grade all values that fall in each grid in the local map according to frequency, showing only the most frequent features as the final category of this grid. Each category such as buildings or cars are discriminated by different number labels. Furthermore, the distance between each grid, denoted as $r$ (5m), is also input to the LLM as the metric information. In this case, the LLM can directly reason about positional relationships among each landmark.
III-C Designing Text Prompt for LLM Planning
Given that LLMs are well-trained in reasoning and planning, we complement their spatial perception and scene understanding capability with formatted prompt and action space. This prompt involves two components. First, the task description is prepared which includes the definition of the environment and the UAV, as well as the format of input and output. We require an LLM to leverage Chain-of-Thought reasoning [wei2022chain] after capturing the observation, thinking step by step in an order of observation-thought-planning-history to predict the next action. Regarding the input instruction, sub-goals are extracted by the LLM, and the landmarks given in the instructions will be updated in the task definition as the scene changes. Second, to improve the quality of prompts, we implemented a cyclic refinement process. This involves designing a new set of instructions for an LLM, including elements such as ‘history action’ and ‘next planning’ corresponding to the STMR module. The task description also emphasizes updating the status of planning, including three states, i.e., todo, in process, and completed. The STMR is gradually updated in this iterative process as the UAV explores to improve the accuracy and relevance of the prompts.
Example of LLM Prompt [Task Description]You are an embodied unmanned aerial vehicle that navigates in the real world. You need to explore between some places marked and ultimately find the destination to stop.
[Input Format]: ‘Instruction’ is a global, step-by-step detailed guidance. ‘History’ is your previously executed actions. ‘Map’ is a top-down coordinate map, using a 20*20 matrix M to represent coordinates. The numbers in the matrix represent different landmarks. You are at [10,10] of the matrix, which is a special value denoting your orientation, e.g., west0 means face due west and the declination angle is 0 degree. The distance between adjacent numbers is 5 meters in the real world. The numbers in the matrix M represent:
[0:Unexplored 1:road 2:building … -1:your past trajectory]. ‘Plan’ records previous long-term multi-step planning info that you can refer to now.
[Output Format] ‘Thought’: your thoughts about the task, which may include your comprehension, surroundings, history, and etc. ‘Observation’: the description of the map. ‘Plan’: your updated plan. ‘Action’: your next action."
Think step by step. First, judge by the ‘Map’, give a first ‘Thought’, and depict your orientation. Second, check that if a landmark in the current ‘plan’ is within 5 meters of your current position, then based on ‘Instruction’ and the previous ‘Plan’, update your multi-step ‘Plan’. Each plan needs to follow a state word (TODO, In Process, Completed). Finally, judge by the ‘Map’ again, and give the specific "Action" according to the action format. Action format: Action: (right, left, lift, down, straight, back), (degree (0-15 degrees)), (distance (0-10 meters)).
  Instruction: lift off and head straight across the water to the road…
History: [lift off 5 meters…]
Map: STMR
Plan:
1. (Completed) Lift off 5 meters.
2. (Completed) Turn 180 degrees.
3. (Completed) Head straight across the water to the road.
4. (In Process) Turn left to align the road.
5. (TODO)…
IV Experiment Setup
TABLE I: Comparison on the validation seen and unseen set of the AerialVLN-S dataset.
Validation Seen Validation Unseen
# AerialVLN-S NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$
S1 Random 109.6 0.0 0.0 149.7 0.0 0.0
S2 Action Sampling 213.8 0.9 5.7 237.6 0.2 1.1
S3 LingUNet [misra2018mapping] 383.8 0.6 6.9 368.4 0.4 3.6
S4 Seq2Seq [anderson2018vision] 146.0 4.8 19.8 218.9 2.3 11.7
S5 CMA [liu_2023_AerialVLN] 121.0 3.0 23.2 172.1 3.2 16.0
S6 LAG [liu_2023_AerialVLN] 90.2 7.2 15.7 127.9 5.1 10.5
S7 Ours 96.3 12.6 31.6 119.5 10.8 23.0
Datasets
We conduct experiments using a novel and complex dataset provided by [liu_2023_AerialVLN], which we refer to as the "AerialVLN-S" dataset, to assess the performance of aerial VLN tasks. This dataset is designed to mimic real-world urban environments with more than 870 different kinds of objects and various scenarios including downtown cities, factories, parks, and villages. Besides, the AerialVLN-S dataset encompasses 8,446 flying paths recorded by experienced human UAV pilots who possess the AOPA (Aircraft Owners and Pilots Association) certificate, ensuring the paths’ realism and quality. Several comparison methods are evaluated on this dataset, including state-of-the-art VLN models, to establish a benchmark for aerial VLN tasks. We utilize AerialVLN-S as the benchmark, with 12 scenes from the validation seen set and 5 scenes from the validation unseen set, respectively. Notably, the test set is not used because it is inaccessible and the evaluation server [liu_2023_AerialVLN] is unavailable now.
Evaluation metrics
To rigorously assess the performance of our method, we utilize a comprehensive set of evaluation metrics as in [liu_2023_AerialVLN]. We focus on several key aspects, i.e., Navigation Error (NE) quantifying the distance between the UAV’s stopping point and the actual destination, Success Rate (SR) which measures the proportion of navigations that successfully reach the destination within a 20-meter threshold, and Oracle Success Rate (OSR), an idealized measure considering any point on the trajectory that comes within 20 meters of the destination as a success.
Implementation details
Our framework is implemented in both a simulator and a real environment. For the simulation environment, we evaluated the algorithm on airsim and UE4, on a laptop with an Intel i9 12th generation CPU and an Nvidia RTX 4090 GPU. For the outdoor environment, we test on a Q250 airframe, carrying an Intel RealSense D435i depth camera, and an NVIDIA Jetson Xavier NX running Ubuntu 18.04 as onboard computer. We use the online API of GPT-4V and GPT-4o, where the default parameters are employed.
IV-A Experimental Results
Quantitative results in simulator
Table I compares our method with several other works. Random and Action Sampling respectively mean sampling actions randomly and according to the action distribution of the training set. LingUNet [misra2018mapping] assumes an agent can see the destination from the beginning, which is transformed into a step-wise paradigm by [liu_2023_AerialVLN]. Seq2Seq [anderson2018vision] is a learning-based recurrent policy, directly predicting the next action. CMA is a baseline model in [liu_2023_AerialVLN], considering attention between multi-modalities. LAG is the main method in [liu_2023_AerialVLN], designing a look-ahead guidance to adjust paths. In comparison, our method significantly outperforms others in most metrics, demonstrating the superiority of our method. We found that instead of processing the entire long instruction and visual observations at each time step, the formatted decomposed sub-goal is more conducive to the fine-grained alignment of each action and landmark. Moreover, UAV has a better understanding of its own spatial position and other semantic constraints under the guidance of our STMR, which has obvious advantages in long instruction environments and leads to more robust execution.
Quantitative results in real environment
To test whether our solution works well in the real world, we collect 10 outdoor scenes including street scenes and forests with ground-truth lengths ranging from 50m to 500m. Then, we apply our method on real UAVs and call cloud-hosted LLMs (GPT-4o is used) to navigate in these challenging environments. LLM-based VLN methods MapGPT [chen2024mapgpt] and NavGPT [zhou2024navgpt] are employed and compared in this experiment. From demo snapshot Fig. 4, we can see that our method can align visual and textual landmarks as well as understand commands. Finally, the UAV reaches the destination successfully.  Table II also shows that the proposed method achieves much better results than most recent LLM-based VLN methods.
Figure 4: Qualitative results in real environment. Short-range movements of the same action are merged into a single long-range movement, as shown in the images sequence.
TABLE II: Qualitative results in real environment.
Module SR/$\uparrow$ OSR/$\uparrow$
MapGPT [chen2024mapgpt] 20 20
NavGPT [zhou2024navgpt] 10 20
Ours 40 70
Failure case analysis
In the simulation, we analyze two most common failures caused by incorrect planning or execution. The first typical failure is the misunderstanding of ambiguous instructions. In some instructions, there are a lot of continuous commands without any landmark, such as "turn left, then move right, then go straight", which have no objects for an LLM to refer to. This case will directly leads to the repeated execution of a step. The second typical error is caused by the inaccuracy of the visual perceptor. Although modern perception models show strong ability, they still need to be improved in identifying objects from different views. As a result, the key landmark may not be mapped to the matrix representation, causing the proper action not to be completed. This observation aligns with our real-world experiment, where most errors are from perceptions.
TABLE III: Ablation study on randomly selected 100 scenes from Validation Unseen dataset for different semantic spatial representations as prompts.
# Method Validation Unseen
NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$
1 Topo 203.3 4.9 12.8
2 Metric 165.0 6.1 13.7
3 Ours 87.1 19.4 26.2
IV-B Comprehensive Analysis
We carry out several ablation studies to assess the effectiveness of the proposed framework. All ablation experiments are conducted on the AerialVLN-S task with the validation unseen dataset. Corresponding results are depicted in Table III,  Table IV and Table V.
Explicit spatial information
To demonstrate the spatial representation ability of STMR, we further compare it with other prompting formats as an LLM’s spatial information prompt. In Topo format (row 1 of Table III, we maintain a linguistic-formed map that captures the topological relationships between different nodes. Each node records textual descriptions of visual observations, and the connectivity between nodes is described using textual prompts. For example, "Place 1 is connected with Places 2, 4, 0, 3". In Metric format (row 2 of Table III, we arrange the visual observations from 8 different directions in a clockwise order relative to the agent’s current orientation and concatenate them into a single prompt. Each vision observation records the direction and distance of the landmark, for example, "a white building in the left front 10 meters away". It is shown that row 2 improves SR by 1.2% (absolute) over row 1 but still poorer navigation performance, since for some landmarks such as roads and rivers, UAVs still cannot obtain region-level spatial information based on a certain point distance description. Our STMR combining the semantic, topological, and metric information significantly enhances the UAV’s exploration capability and improves the accuracy in SR and OSR by over 10%.
State update of sub-goal
To ensure that each sub-goal will be executed despite environment changes, in each iteration, the LLM does not change the original path plan, but updates the status of the current sub-goal according to the observation, including todo, in process, and completed. We compare this strategy with updating the entire multi-step path planning (row 1) according to the observation in each iteration. The results are shown in Table IV. This strategy does not notably enhance the Oracle Success Rate (OSR), showing that the proposed STMR already furnishes enough capacity for navigating the environment comprehensively. Nonetheless, it positively impacts the decision-making process, resulting in an increase in the success rate (SR) from 9.0% to 10.8%.
TABLE IV: Ablation study for the state update.
Module NE/$\downarrow$ SR/$\uparrow$ OSR/$\uparrow$
w/o state 289.7 9.0 22.1
w state 219.8 10.8 23.0
TABLE V: Ablation study for the direct use of RGB images and the proposed STMR as the visual prompt for LLMs.
LLM Input RGB STMR NE/$\downarrow$ SR/$\uparrow$ OSR/$\uparrow$
GPT-4V $\checkmark$ 350.0 1.5 9.7
$\checkmark$ 193.1 5.6 17.4
GPT-4o $\checkmark$ 412.0 1.1 10.0
$\checkmark$ 119.5 10.8 23.0
Direct visual input and STMR
Due to the strong multi-modal capabilities of GPT-4V and GPT-4o, they can reason with both visual and textual prompts. However, it is challenging for them to understand landmark semantics directly from step-by-step input RGB images. This may be because the inability of LLMs in spatial-aware reasoning. As shown in Table V, the proposed STMR representation with LLMs significantly outperforms the direct input of RGB images with VLMs.
V CONCLUSIONS AND FUTURE WORK
This paper addresses the challenging aerial VLN task by introducing an LLM-based end-to-end framework. To enhance the spatial reasoning ability of LLMs, we propose Semantic-Topo-Metric Representation (STMR). STMR first integrates instruction-related landmarks and their locations into a top-down map, and subsequently transforms this map into a matrix representation containing semantic, topological, and distance metric information. Taking the proposed STMR as a part of prompts of the LLM, we significantly improve the UAV’s navigation capabilities. Our framework achieves state-of-the-art results on the AerialVLN-S dataset, demonstrating its effectiveness. In the future, we will explore flexible framework that will interact with humans to eliminate ambiguities in the instructions.

### v2

@@
-Aerial Vision-and-Language Navigation via Semantic-Topo-Metric Representation Guided LLM Reasoning
+Exploring Spatial Representation to Enhance LLM Reasoning in Aerial Vision-Language Navigation
 Abstract
 Aerial Vision-and-Language Navigation (VLN) is a novel task enabling Unmanned Aerial Vehicles (UAVs) to navigate in outdoor environments through natural language instructions and visual cues.
-It remains challenging due to the complex spatial relationships in outdoor aerial scenes.
-In this paper, we propose an end-to-end zero-shot framework for aerial VLN tasks, where the large language model (LLM) is introduced as our agent for action prediction.
-Specifically, we develop a novel Semantic-Topo-Metric Representation (STMR) to enhance the spatial reasoning ability of LLMs.
-This is achieved by extracting and projecting instruction-related semantic masks of landmarks into a top-down map that contains the location information of surrounding landmarks.
-Further, this map is transformed into a matrix representation with distance metrics as the text prompt to the LLM, for action prediction according to the instruction.
-Experiments conducted in real and simulation environments have successfully proved the effectiveness and robustness of our method, achieving 15.9% and 12.5% improvements (absolute) in Oracle Success Rate (OSR) on AerialVLN-S dataset.
+However, it remains challenging due to the complex spatial relationships in outdoor aerial scenes.
+In this paper, we propose a training-free, zero-shot framework for aerial VLN tasks, where the large language model (LLM) is leveraged as the agent for action prediction.
+Specifically, we develop a novel Semantic-Topo-Metric Representation (STMR) to enhance the spatial reasoning capabilities of LLMs.
+This is achieved by extracting and projecting instruction-related semantic masks of landmarks onto a top-down map, which contains location information about surrounding landmarks.
+The map is then transformed into a matrix representation with distance metrics, serving as the text prompt to the LLM for action prediction in response to the given instruction.
+Experiments conducted in real and simulation environments have successfully proved the effectiveness and robustness of our method, achieving at most 10.8% improvements (absolute) in Oracle Success Rate (OSR) on AerialVLN-S and Openfly dataset.
 I Introduction
@@
 Figure 1: The pipeline to obtain STMR.
-Observation image, semantic mask, and depth image (a) will be aggregated in the top-down map (red dashed box) (b), which captures the Semantic-Topo-Metric information of the neighborhood of the current position, further generating a 20x20 matrix representation (c).
+Observation image, semantic mask, and depth image (a) can be aggregated in the top-down map (red dashed box) (b), which captures the Semantic-Topo-Metric information of the neighborhood of the current position, further generating a 20x20 matrix representation (c) for LLM reasoning.
 Although existing VLN methods have made significant progress, most of them are designed for indoor or ground-based outdoor environments and pay less attention to aerial VLN tasks.
@@
 To overcome these challenges, we propose a zero-shot LLM-based aerial VLN framework, which takes natural language instructions, RGB images, and depth images as input, and directly makes action predictions (e.g., forward 10 meters) through single-step reasoning updates.
-Specifically, a Semantic-Topo-Metric Representation (STMR) is designed for LLMs prompting.
-We first extract instruction-related landmarks and obtain corresponding semantic masks by perception models, i.e., Grounding Dino [liu2023grounding] and Tokenize Anything [pan2023tap].
+Specifically, a Semantic-Topo-Metric Representation (STMR) is designed for LLMs spatial reasoning.
+Leveraging the developments in open-vocabulary perception models [zhang2018review, jiang2018deep, minderer2022simple, cheng2024yolo], we first extract instruction-related landmarks and obtain corresponding semantic masks by perception models, i.e., Grounding Dino [liu2023grounding] and Tokenize Anything [pan2023tap].
 After that, semantic masks obtained from each step are gradually projected into a top-down map using depth and pose transformation as shown in Fig.
@@
 Our contributions are summarized as follows:
-  • To our knowledge, we propose the first LLM-based end-to-end framework for the aerial VLN task, facilitating the development of UAV navigators.
-This framework requires no training and no extra low-level action planner, thus allowing for easy integration.
+  • To our knowledge, we design the first LLM-based training-free framework for the aerial VLN task, facilitating the development of UAV navigators.
+Without extra action planner, the proposed framework allows for easy integration.
   • We propose the Semantic-Topo-Metric Representation (STMR), a unique matrix representation that encompasses topological, semantic, and metric information.
-STMR is designed to significantly enhance the spatial-aware reasoning capabilities of LLMs in outdoor environments.
-  • Extensive experiments on the aerial VLN task especially in real environments demonstrate that the proposed method outperforms previous state-of-the-art methods by a large margin, establishing a strong baseline for future zero-shot Aerial VLN tasks.
+STMR is designed to enhance the spatial-aware reasoning capabilities of LLMs in outdoor environments.
+  • Extensive experiments on the aerial VLN task demonstrate that the proposed method outperforms previous state-of-the-art methods by a large margin, establishing a strong baseline for future zero-shot Aerial VLN tasks.
 II Related Work
@@
 More recent works [zhou2024navgpt, chen2024mapgpt] leverage the reasoning and dialogue capabilities of LLMs, achieving great progress.
-However, most of them are developed under a ground-based, discrete VLN setting, which simplifies navigation as traversing on a predefined graph on the ground.
-This limits the free movement space of UAVs in the real world.
+However, most of them are developed under a ground-based, discrete VLN setting, which simplifies navigation by traversing on a predefined graph on the ground.
+This limits the UAVs’ free movement space in the real world.
 Aerial VLN remains challenging due to the large-scale and complex environments.
@@
 [misra2018mapping] decompose instruction execution in two stages, separately using supervised learning for goal prediction and policy gradient for action generation.
-AerialVLN [liu_2023_AerialVLN] contributes a much more challenging aerial VLN dataset focusing on large-scale environments, and provides a look-ahead guidance method as the baseline.
+AerialVLN [liu_2023_AerialVLN] contributes a much more challenging aerial VLN dataset focusing on outdoor aerial environments, and provides a look-ahead guidance method as the baseline.
+OpenFly [openfly] constructs a large-scale dataset across several high-quality scenes, and proposes a keyframe-aware VLN model.
 Despite the progress, the generalizability and performance of these methods still require improvement.
@@
 To achieve the navigation goal, the UAV considers both the instruction and visual perceptions, and predicts an action from the action space (right, left, up, down, forward, backward) with corresponding parameters at each time step $t$.
-Navigation ends when the UAV predicts a ’Stop’ action or reaches a pre-defined maximum action number.
-The navigation is recognized as a success if the UAV stops at a location that is less than 20 meters from the target location.
+Navigation ends when the UAV predicts a ‘Stop’ action or reaches a pre-defined maximum action number.
+Following [liu_2023_AerialVLN], the navigation is recognized as a success if the UAV stops at a location that is less than 20 meters from the target location.
 Figure 2: Our method consists of three modules, i.e.
-sub-goal extracting, Semantic-Topo-Metric Representating, and LLM planner.
+sub-goal extracting, Semantic-Topo-Metric Representation, and LLM planner.
 First, we prompt an LLM to break down long instructions and extract specific landmarks, which will be classified by perception models, generating corresponding semantic masks.
@@
 Previous LLM-based VLN methods use natural language to describe current observations, or nodes and edges of a topological graph to model the spatial information of the environment.
-However, in open scenarios, simple directional words such as "next to" or "aside" are not enough to describe complex spatial relationships, which can easily introduce ambiguity into LLMs.
-To address this challenge, we introduce the Semantic-Topo-Metric Representation (STMR) to enhance the spatial-aware reasoning capability of LLMs.
+However, in open scenarios, simple directional words such as “next to” or “aside” are not enough to describe complex spatial relationships, which can easily introduce ambiguity into LLMs.
+To address this challenge, we introduce the STMR to enhance the spatial-aware reasoning capability of LLMs.
 Specifically, STMR incrementally takes an RGB image $I_{t}^{R}$ and a depth map $I_{t}^{D}$ as input from each step, and generates a dynamically updated matrix representation with semantic, topological, and metric information as its output.
 The details of STMR are presented as follows.
-2D Visual Perception
-Impressed by the powerful open-vocabulary detection capabilities of Grounding DINO $\operatorname{GD}(\cdot)$, as well as the captioning and segmentation capabilities of Tokenize Anything model $\operatorname{TA}(\cdot)$, we integrate these two models as our 2D visual perceptor, as illustrated in Fig.
+III-B1 2D Visual Perception
+Impressed by the powerful open-vocabulary detection capabilities of Grounding DINO $\operatorname{GD}(\cdot)$, as well as the captioning and segmentation capabilities of Tokenize Anything model $\operatorname{TA}(\cdot)$, we integrate these two models as our 2D visual preceptor, as illustrated in Fig.
 3.
-Given a single RGB image $I_{t}^{R}$ and an instruction $L$ as input, we first obtain detailed landmark categories using a Landmark Extractor $\operatorname{LE}(\cdot)$ driven by an LLM, and then identify the corresponding bounding box for each category through $\operatorname{GD}(\cdot)$.
-Next, we employ $\operatorname{TA}(\cdot)$ to take each bounding box as a prompt and output a set of 2D semantic masks $m^{(t)}$ and captions $z^{(t)}$ for the current RGB image $I_{t}^{R}$.
+Given a single RGB image $I_{t}^{R}$ and an instruction $L$ as input, we first obtain detailed landmark categories $C=\{c_{1},c_{2},...,c_{n}\}$ using a Landmark Extractor $\operatorname{LE}(\cdot)$ driven by an LLM, and then identify the corresponding bounding box for each category through $\operatorname{GD}(\cdot)$.
+Next, we employ $\operatorname{TA}(\cdot)$ to take each bounding box as a prompt and output a set of 2D semantic masks $m^{(t)}$ and captions $h^{(t)}$ for the current RGB image $I_{t}^{R}$.
 The entire process can be described as:
 Figure 3: 2D Visual Perceptor for the UAV.
-$\{m^{(t)},\ z^{(t)}\}=\operatorname{TA}\left({I_{t}^{R}},\ \operatorname{GD}\left({I_{t}^{R}},\ \operatorname{LE}({L})\right)\right).$ (1)
-To improve the simplicity of the semantic masks and reduce the misleading of LLM-reasoning caused by numerous open-vocabulary categories, we propose a text-matching method to mitigate over-classification.
+$\{m^{(t)},\ h^{(t)}\}=\operatorname{TA}\left({I_{t}^{R}},\ \operatorname{GD}\left({I_{t}^{R}},\ \operatorname{LE}({L})\right)\right).\$ (1)
+To improve the simplicity of the semantic masks and reduce the misleading of LLM reasoning caused by numerous open-vocabulary categories, we propose a text-matching method to mitigate over-classification.
 As shown in Fig.
-3, our method involves vectorizing the landmarks extracted from the instruction and the captions $z^{(t)}$ generated in each $I_{t}^{R}$.
+3, our method involves vectorizing the landmarks extracted from the instruction and the captions $h^{(t)}$ generated in each $I_{t}^{R}$.
 Then, we calculate the cosine similarity between these vectorized landmarks and captions using TF-IDF [qaiser2018text].
-If the similarity score exceeds the threshold $\tau$ >0.8, the landmark is classified as visible in the current view.
+If the similarity score exceeds the threshold $\tau$ (0.8), the landmark is classified as visible in the current view, and other irrelevant masks will be ignored.
 By implementing this strategy, we effectively simplify semantic masks and ensure the LLM’s reasoning focuses on relevant categories.
-Sub-goal-driven top-down map
+III-B2 Sub-goal-driven top-down map
 Considering that the top-down view can better express spatial relationships than the first-person view, we further utilize the depth image $I_{t}^{D}$ to map semantic masks to 3D space, and then convert them into a top-down map.
-Specifically, with the RGB image $I_{t}^{R}$ processed by the 2D visual perceptor, we produce first-person view predictions by assigning semantic labels to each pixel, resulting in a segmented image with identified objects and regions.
-Subsequently, the depth image $I_{t}^{D}$ is converted into a 3D point cloud, where each pixel is mapped to a 3D point $(X,Y,Z)$ based on its depth value and camera parameters:
-$X=\frac{(u-c_{x})I_{t}^{D}(u,v)}{f_{x}},Y=\frac{(v-c_{y})I_{t}^{D}(u,v)}{f_{y}},Z=I_{t}^{D}(u,v),$ (2)
+Specifically, by applying the above 2D visual perception process to the RGB image $I_{t}^{R}$, we obtain a segmented image with clearly identified objects and regions.
+Subsequently, the depth image $I_{t}^{D}$ is converted into a 3D point cloud, where each pixel is mapped to a corresponding 3D point $(X,Y,Z)$ based on its depth value and camera parameters:
+$X=\frac{(u-c_{x})I_{t}^{D}(u,v)}{f_{x}},Y=\frac{(v-c_{y})I_{t}^{D}(u,v)}{f_{y}},$ (2)
+$Z=I_{t}^{D}(u,v),$ (3)
 where $(u,v)$ are the pixel coordinates, $(c_{x},c_{y})$ are the camera’s principal point coordinates, and $(f_{x},f_{y})$ are the focal lengths.
@@
 At each step, the position of a UAV will also be recorded as waypoints with special values and projected onto the top-down map.
-Matrix representation
+III-B3 Matrix representation
 We found that directly inputting the image-format top-down map to a vision-language model (VLM) often yields poor reasoning results.
-Therefore, we process the visual map into a text-based matrix representation, which is input to an LLM to guide the navigation.
-Specifically, the UAV incrementally collects observations and gradually constructs the sub-goal-driven top-down map of the surrounding environment.
+Alternatively, we process the visual map into a text-based matrix representation, which is input to an LLM to guide the navigation.
+Specifically, the UAV incrementally collects observations and gradually constructs the top-down map of the surrounding environment.
 Considering the perception quality of onboard sensors and the effective processing capacity of modern LLMs, we define a local map centered at the UAV’s current position as shown in Fig.
@@
 Particularly, we grade all values that fall in each grid in the local map according to frequency, showing only the most frequent features as the final category of this grid.
-Each category such as buildings or cars are discriminated by different number labels.
-Furthermore, the distance between each grid, denoted as $r$ (5m), is also input to the LLM as the metric information.
+Each category such as building or car is discriminated by different number labels.
+Furthermore, the grid size $r$ (5m) is also input to the LLM as the metric information (see Table VII for ablation study on the grid size).
 In this case, the LLM can directly reason about positional relationships among each landmark.
@@
 This involves designing a new set of instructions for an LLM, including elements such as ‘history action’ and ‘next planning’ corresponding to the STMR module.
-The task description also emphasizes updating the status of planning, including three states, i.e., todo, in process, and completed.
+The task description also emphasizes updating the status of planning, including three states, i.e., todo, in process, and completed, in order to minimize the problem of hallucinations that is common in LLM.
 The STMR is gradually updated in this iterative process as the UAV explores to improve the accuracy and relevance of the prompts.
-Example of LLM Prompt [Task Description]You are an embodied unmanned aerial vehicle that navigates in the real world.
+Example of LLM Prompt (simplified) [Task Description]You are an embodied unmanned aerial vehicle that navigates in the real world.
 You need to explore between some places marked and ultimately find the destination to stop.
-[Input Format]: ‘Instruction’ is a global, step-by-step detailed guidance.
+[Input Format] ‘Instruction’ is a global, step-by-step detailed guidance.
 ‘History’ is your previously executed actions.
-‘Map’ is a top-down coordinate map, using a 20*20 matrix M to represent coordinates.
-The numbers in the matrix represent different landmarks.
-You are at [10,10] of the matrix, which is a special value denoting your orientation, e.g., west0 means face due west and the declination angle is 0 degree.
+‘Map’ is a top-down coordinate map, using a 20*20 grid M to represent coordinates, heading towards upper of the map.
+The value of the grid represent different landmarks.
+You are at [9,9] on the ’Map’ (marked as P).
 The distance between adjacent numbers is 5 meters in the real world.
 The numbers in the matrix M represent:
-[0:Unexplored 1:road 2:building … -1:your past trajectory]. ‘Plan’ records previous long-term multi-step planning info that you can refer to now.
+[0:Unexplored 1:your past trajectory 2:road 3:building … ]. ‘Plan’ records previous long-term multi-step planning info that you can refer to now.
 [Output Format] ‘Thought’: your thoughts about the task, which may include your comprehension, surroundings, history, and etc.
@@
 ‘Plan’: your updated plan.
-‘Action’: your next action."
+‘Action’: your next action.”
 Think step by step.
-First, judge by the ‘Map’, give a first ‘Thought’, and depict your orientation.
-Second, check that if a landmark in the current ‘plan’ is within 5 meters of your current position, then based on ‘Instruction’ and the previous ‘Plan’, update your multi-step ‘Plan’.
+First, describe your orientation and the landmark in your left, right and front.
+Second, according the ’map’, update your multi-step ‘Plan’.
 Each plan needs to follow a state word (TODO, In Process, Completed).
-Finally, judge by the ‘Map’ again, and give the specific "Action" according to the action format.
-Action format: Action: (right, left, lift, down, straight, back), (degree (0-15 degrees)), (distance (0-10 meters)).
-  Instruction: lift off and head straight across the water to the road…
+Finally, give the specific “Action” according to the action format: Action: (right, left, lift, down, straight, back), (degree (0-90 degrees)), (distance (between 0-10 meters)).
+  A Simplified Input in the Cyclic Process
+Instruction
+: lift off and head straight across the water to the road…
 History: [lift off 5 meters…]
@@
 Plan:
-1. (Completed) Lift off 5 meters.
-2. (Completed) Turn 180 degrees.
-3. (Completed) Head straight across the water to the road.
-4. (In Process) Turn left to align the road.
+1. Lift off 5 meters. (Completed)
+2. Turn 180 degrees. (Completed)
+3. Head straight across the water to the road. (Completed)
+4. Turn left to align the road. (In Process)
 5. (TODO)…
-IV Experiment Setup
-TABLE I: Comparison on the validation seen and unseen set of the AerialVLN-S dataset.
-Validation Seen Validation Unseen
-# AerialVLN-S NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$
-S1 Random 109.6 0.0 0.0 149.7 0.0 0.0
-S2 Action Sampling 213.8 0.9 5.7 237.6 0.2 1.1
-S3 LingUNet [misra2018mapping] 383.8 0.6 6.9 368.4 0.4 3.6
-S4 Seq2Seq [anderson2018vision] 146.0 4.8 19.8 218.9 2.3 11.7
-S5 CMA [liu_2023_AerialVLN] 121.0 3.0 23.2 172.1 3.2 16.0
-S6 LAG [liu_2023_AerialVLN] 90.2 7.2 15.7 127.9 5.1 10.5
-S7 Ours 96.3 12.6 31.6 119.5 10.8 23.0
-Datasets
-We conduct experiments using a novel and complex dataset provided by [liu_2023_AerialVLN], which we refer to as the "AerialVLN-S" dataset, to assess the performance of aerial VLN tasks.
-This dataset is designed to mimic real-world urban environments with more than 870 different kinds of objects and various scenarios including downtown cities, factories, parks, and villages.
-Besides, the AerialVLN-S dataset encompasses 8,446 flying paths recorded by experienced human UAV pilots who possess the AOPA (Aircraft Owners and Pilots Association) certificate, ensuring the paths’ realism and quality.
-Several comparison methods are evaluated on this dataset, including state-of-the-art VLN models, to establish a benchmark for aerial VLN tasks.
-We utilize AerialVLN-S as the benchmark, with 12 scenes from the validation seen set and 5 scenes from the validation unseen set, respectively.
-Notably, the test set is not used because it is inaccessible and the evaluation server [liu_2023_AerialVLN] is unavailable now.
-Evaluation metrics
-To rigorously assess the performance of our method, we utilize a comprehensive set of evaluation metrics as in [liu_2023_AerialVLN].
-We focus on several key aspects, i.e., Navigation Error (NE) quantifying the distance between the UAV’s stopping point and the actual destination, Success Rate (SR) which measures the proportion of navigations that successfully reach the destination within a 20-meter threshold, and Oracle Success Rate (OSR), an idealized measure considering any point on the trajectory that comes within 20 meters of the destination as a success.
-Implementation details
+IV Experiment
+TABLE I: Comparison on the validation seen of the AerialVLN-S. * denotes we train these methods by running their source codes.
+Validation Seen
+model NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ SDTW/$\uparrow$
+Random 109.6 0.0 0.0 0.0
+Action Sampling 213.8 0.9 5.7 0.3
+LingUNet [misra2018mapping] 383.8 0.6 6.9 0.2
+Seq2Seq [anderson2018vision] 146.0 4.8 19.8 1.6
+CMA [CMA] 121.0 3.0 23.2 0.6
+AerialVLN [liu_2023_AerialVLN] 90.2 7.2 15.7 2.4
+Navid* [navid] 105.1 6.8 15.5 1.1
+Openfly* [openfly] 127.2 8.1 21.8 1.6
+Ours 96.3 12.6 31.6 2.7
+TABLE II: Comparison on the validation unseen of the AerialVLN-S. * denotes we train these methods by running their source codes.
+Validation Unseen
+model NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ SDTW/$\uparrow$
+Random 149.7 0.0 0.0 0.0
+Action Sampling 237.6 0.2 1.1 0.1
+LingUNet [misra2018mapping] 368.4 0.4 3.6 0.9
+Seq2Seq [anderson2018vision] 218.9 2.3 11.7 0.7
+CMA [CMA] 172.1 3.2 16.0 1.1
+AerialVLN [liu_2023_AerialVLN] 127.9 5.1 10.5 1.4
+Navid* [navid] 106.9 6.1 12.3 0.7
+Openfly* [openfly] 113.8 7.6 18.2 1.5
+Ours 119.5 10.8 23.0 1.9
+TABLE III: Comparison on OpenFly dataset unseen split.
+OpenFly Unseen
+model NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ SDTW/$\uparrow$
+Random 301m 0.1 0.1 0
+Seq2Seq[anderson2018vision] 291m 0.1 3.2 0.1
+CMA[CMA] 268m 0.7 6.6 0.1
+AerialVLN[liu_2023_AerialVLN] 227m 2.3 19.8 1.3
+Navid[navid] 200m 2.6 16.9 2.2
+OpenFly[openfly] 198m 5.1 27.3 2.0
+Ours 205m 8.0 31.0 2.4
+IV-A Dataset and Implementation Details
+IV-A1 Dataset
+We conduct experiments using two datasets to assess aerial VLN task performance, i.e., AerialVLN-S benchmark from [liu_2023_AerialVLN] and the recently introduced OpenFly dataset [openfly].
+The AerialVLN-S dataset simulates real-world urban environments with over 870 object categories across diverse scenarios such as downtown cities, factories, parks, and villages.
+It contains 8,446 expert-curated flying paths recorded by AOPA-certified UAV pilots, ensuring trajectory realism and navigational fidelity.
+For evaluation, we utilize 12 scenes from its validation set and 5 scenes from the validation unseen set, as the official test set remains inaccessible.
+To enable broader comparative analysis, we introduce the OpenFly unseen dataset, which provides 1,200 test trajectories uniformly sampled from more diverse scenarios.
+We compare with several state-of-the-art VLN models on both benchmarks, using OpenFly’s test unseen trajectories for generalization assessment.
+This dual evaluation strategy rigorously tests the performances of these works, establishing comprehensive baselines for aerial VLN research.
+IV-A2 Evaluation metrics
+To validate the effectiveness of our method, we utilize a comprehensive set of evaluation metrics following [liu_2023_AerialVLN].
+We focus on several key aspects, i.e., Navigation Error (NE) quantifying the distance between the UAV’s stopping point and the actual destination, Success Rate (SR) which measures the proportion of navigations that successfully reach the destination within a 20-meter threshold, Oracle Success Rate (OSR), an idealized measure considering any point on the trajectory that comes within 20 meters of the destination as a success, and Success rate weighted by Normalised Dynamic Time Warping (SDTW), which takes into account both the navigation success rate and the similarity between the real paths.
+IV-A3 Implementation details
 Our framework is implemented in both a simulator and a real environment.
-For the simulation environment, we evaluated the algorithm on airsim and UE4, on a laptop with an Intel i9 12th generation CPU and an Nvidia RTX 4090 GPU.
-For the outdoor environment, we test on a Q250 airframe, carrying an Intel RealSense D435i depth camera, and an NVIDIA Jetson Xavier NX running Ubuntu 18.04 as onboard computer.
-We use the online API of GPT-4V and GPT-4o, where the default parameters are employed.
-IV-A Experimental Results
-Quantitative results in simulator
-Table I compares our method with several other works.
+In the simulation environments of AirSim and Unreal Engine 4 (UE4), we evaluated the methods using a PC equipped with an Intel i9 12th-generation CPU and Nvidia RTX 4090 GPUs.
+For the real world environment, we test on a Q250 airframe, carrying an Intel RealSense D435i depth camera, and an NVIDIA Jetson Xavier NX running Ubuntu 18.04 as onboard computer.
+All methods are run on an external PC (same as the simulation setup) communicating with the onboard computer via a UDP network.
+For LLM reasoning we use the online API of GPT-4V and GPT-4o, where the default parameters are employed.
+IV-B Experimental Results
+IV-B1 Quantitative results in simulator
+Table I and  Table II compare our method with several other works on the AerialVLN-S dataset.
 Random and Action Sampling respectively mean sampling actions randomly and according to the action distribution of the training set.
@@
 LAG is the main method in [liu_2023_AerialVLN], designing a look-ahead guidance to adjust paths.
+Navid [navid] is a VLM-based method, achieving great success in the indoor VLN task.
+OpenFly [openfly] proposes a keyframe-aware aerial VLN model, taking key historical observations into account.
 In comparison, our method significantly outperforms others in most metrics, demonstrating the superiority of our method.
+Notably, our method is training-free, while others are trained using corresponding data.
+Table III shows the generalization results of different methods on the OpenFly Unseen dataset, where our method also achieves better performance than other works.
 We found that instead of processing the entire long instruction and visual observations at each time step, the formatted decomposed sub-goal is more conducive to the fine-grained alignment of each action and landmark.
 Moreover, UAV has a better understanding of its own spatial position and other semantic constraints under the guidance of our STMR, which has obvious advantages in long instruction environments and leads to more robust execution.
-Quantitative results in real environment
+Aerial VLN is a challenging task that has not been fully studied, and the overall performance needs to be improved.
+We believe that our work can bring some insights into this field.
+IV-B2 Quantitative results in real environment
 To test whether our solution works well in the real world, we collect 10 outdoor scenes including street scenes and forests with ground-truth lengths ranging from 50m to 500m.
-Then, we apply our method on real UAVs and call cloud-hosted LLMs (GPT-4o is used) to navigate in these challenging environments.
-LLM-based VLN methods MapGPT [chen2024mapgpt] and NavGPT [zhou2024navgpt] are employed and compared in this experiment.
-From demo snapshot Fig.
-4, we can see that our method can align visual and textual landmarks as well as understand commands.
+Then, we evaluate our method using a real UAV and call cloud-hosted LLMs to navigate in these challenging environments.
+Not only pretrained VLN models but also other LLM-based VLN methods, i.e., MapGPT [chen2024mapgpt] and NavGPT [zhou2024navgpt], are employed and compared in this experiment.
+Notably, MapGPT and NavGPT rely on predefined topological maps to predict the next navigation waypoint, we manually define candidate waypoints for them in this experiment.
+From demonstration snapshot Fig.
+6, we can see that our method can align visual and textual landmarks as well as understand commands.
 Finally, the UAV reaches the destination successfully.
-Table II also shows that the proposed method achieves much better results than most recent LLM-based VLN methods.
-Figure 4: Qualitative results in real environment. Short-range movements of the same action are merged into a single long-range movement, as shown in the images sequence.
-TABLE II: Qualitative results in real environment.
-Module SR/$\uparrow$ OSR/$\uparrow$
-MapGPT [chen2024mapgpt] 20 20
-NavGPT [zhou2024navgpt] 10 20
-Ours 40 70
-Failure case analysis
-In the simulation, we analyze two most common failures caused by incorrect planning or execution.
+Table IV also shows that the proposed method achieves a much better performance than both pretrained VLN methods and training-free LLM-based methods.
+However, we can see that LLM-based methods including ours need more computation time.
+This is mainly because of the latency of the LLM API.
+We believe that with the development of LLMs and hardware computing power, the latency will gradually decrease.
+Figure 4: Demonstration of the STMR in spatial reasoning.
+Figure 5: Failure cases of our method.
+TABLE IV: Quantitative results in real environment. The last column shows the average runtime of aerial VLN procedures.
+model SR/$\uparrow$ OSR/$\uparrow$ Time (s)
+AerialVLN [liu_2023_AerialVLN] 0 10 82
+Navid [navid] 10 10 75
+Openfly [openfly] 20 30 63
+MapGPT [chen2024mapgpt] 20 20 176
+NavGPT [zhou2024navgpt] 10 20 150
+Ours 40 70 117
+Figure 6: Qualitative results in real environment. Short-range movements of the same action are merged into a single long-range movement, as shown in the image sequence.
+IV-B3 Case analysis
+Fig.
+4 demonstrates the effectiveness of our STMR in helping spatial reasoning, where the left part shows an intermediate STMR during the navigation process, and the right part shows simplified prompt, questions, and reasoning answers.
+From Fig.
+4 we can see that an LLM can understand and describe the surrounding environment through the STMR.
+It can even make accurate judgments about objects or areas at a distance.
+Such an ability is extremely helpful for aerial VLN in vast outdoor environments.
+We also summarize two most common failures caused by incorrect planning or execution.
 The first typical failure is the misunderstanding of ambiguous instructions.
-In some instructions, there are a lot of continuous commands without any landmark, such as "turn left, then move right, then go straight", which have no objects for an LLM to refer to.
-This case will directly leads to the repeated execution of a step.
-The second typical error is caused by the inaccuracy of the visual perceptor.
-Although modern perception models show strong ability, they still need to be improved in identifying objects from different views.
-As a result, the key landmark may not be mapped to the matrix representation, causing the proper action not to be completed.
-This observation aligns with our real-world experiment, where most errors are from perceptions.
-TABLE III: Ablation study on randomly selected 100 scenes from Validation Unseen dataset for different semantic spatial representations as prompts.
-# Method Validation Unseen
-NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$
-1 Topo 203.3 4.9 12.8
-2 Metric 165.0 6.1 13.7
-3 Ours 87.1 19.4 26.2
-IV-B Comprehensive Analysis
-We carry out several ablation studies to assess the effectiveness of the proposed framework.
-All ablation experiments are conducted on the AerialVLN-S task with the validation unseen dataset.
-Corresponding results are depicted in Table III,  Table IV and Table V.
-Explicit spatial information
+For example, as shown in the left part of Fig.
+5, sometimes there are a lot of continuous commands without any landmark, such as “turn left, then move right, then go straight”, which have no objects for an LLM to refer to.
+This case will directly lead to the repeated execution of an action.
+The second typical error is caused by the inaccuracy of the visual perception.
+Although modern perception models show strong capability, they still need to be improved in identifying objects from different views.
+As a result, the key landmark may not be mapped to the STMR, causing the proper action not to be completed as illustrated in the right part of Fig.
+5.
+This observation aligns with our real-world experiment.
+TABLE V: Ablation study on different spatial representations.
+Method Validation Unseen
+NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ SDTW/$\uparrow$
+Topo 203.3 4.9 12.8 1.8
+Metric 165.0 6.1 13.7 2.2
+Ours 87.1 19.4 26.2 5.7
+IV-C Ablation Studies
+We carry out comprehensive ablation studies to assess the core designs of the proposed method.
+All ablation experiments are conducted on the randomly selected 100 samples from the AerialVLN-S dataset.
+Corresponding results are depicted in Table V,  Table VI,  Table VII, and Table VIII.
+IV-C1 Different types of spatial information
 To demonstrate the spatial representation ability of STMR, we further compare it with other prompting formats as an LLM’s spatial information prompt.
-In Topo format (row 1 of Table III, we maintain a linguistic-formed map that captures the topological relationships between different nodes.
+In Topo format (row 1 of Table V), we maintain a linguistically formed map that captures the topological relationships between different nodes.
 Each node records textual descriptions of visual observations, and the connectivity between nodes is described using textual prompts.
-For example, "Place 1 is connected with Places 2, 4, 0, 3".
-In Metric format (row 2 of Table III, we arrange the visual observations from 8 different directions in a clockwise order relative to the agent’s current orientation and concatenate them into a single prompt.
-Each vision observation records the direction and distance of the landmark, for example, "a white building in the left front 10 meters away".
-It is shown that row 2 improves SR by 1.2% (absolute) over row 1 but still poorer navigation performance, since for some landmarks such as roads and rivers, UAVs still cannot obtain region-level spatial information based on a certain point distance description.
+For example, “Place 1 is connected with Places 2, 4, 0, 3”.
+In Metric format (row 2 of Table V), we arrange the visual observations from 8 different directions in a clockwise order relative to the agent’s current orientation and concatenate them into a single prompt.
+Each vision observation records the direction and distance of the landmark, for example, “a white building in the left front 10 meters away”.
+It is shown that row 2 improves SR by 1.2% (absolute) over row 1 but still has poorer navigation performance, since for some landmarks such as roads and rivers, UAVs still cannot obtain region-level spatial information based on a certain point distance description.
 Our STMR combining the semantic, topological, and metric information significantly enhances the UAV’s exploration capability and improves the accuracy in SR and OSR by over 10%.
-State update of sub-goal
-To ensure that each sub-goal will be executed despite environment changes, in each iteration, the LLM does not change the original path plan, but updates the status of the current sub-goal according to the observation, including todo, in process, and completed.
-We compare this strategy with updating the entire multi-step path planning (row 1) according to the observation in each iteration.
-The results are shown in Table IV.
+IV-C2 Direct visual input vs STMR
+Due to the powerful multimodal representation capabilities of GPT-4V and GPT-4o, they are capable of direct reasoning with visual and textual prompts.
+However, it is challenging for them to make correct VLN decisions from sequential images.
+This may be because of the inability of VLMs in spatial-aware reasoning.
+As shown in Table VI, the proposed STMR representation with LLMs significantly outperforms the direct input of RGB images into VLMs.
+IV-C3 Grid size of STMR
+[…]
+IV-C4 State update of sub-goal
+To ensure that each sub-goal can be executed despite environmental changes, in each iteration, the LLM does not change the original path plan, but updates the status of the current sub-goal including todo, in process, and completed.
+We compare this strategy with updating the entire multi-step path planning (row 1 of Table VIII) according to the observation in each iteration.
 This strategy does not notably enhance the Oracle Success Rate (OSR), showing that the proposed STMR already furnishes enough capacity for navigating the environment comprehensively.
 Nonetheless, it positively impacts the decision-making process, resulting in an increase in the success rate (SR) from 9.0% to 10.8%.
-TABLE IV: Ablation study for the state update.
-Module NE/$\downarrow$ SR/$\uparrow$ OSR/$\uparrow$
-w/o state 289.7 9.0 22.1
-w state 219.8 10.8 23.0
-TABLE V: Ablation study for the direct use of RGB images and the proposed STMR as the visual prompt for LLMs.
-LLM Input RGB STMR NE/$\downarrow$ SR/$\uparrow$ OSR/$\uparrow$
-GPT-4V $\checkmark$ 350.0 1.5 9.7
-$\checkmark$ 193.1 5.6 17.4
-GPT-4o $\checkmark$ 412.0 1.1 10.0
-$\checkmark$ 119.5 10.8 23.0
-Direct visual input and STMR
-Due to the strong multi-modal capabilities of GPT-4V and GPT-4o, they can reason with both visual and textual prompts.
-However, it is challenging for them to understand landmark semantics directly from step-by-step input RGB images.
-This may be because the inability of LLMs in spatial-aware reasoning.
-As shown in Table V, the proposed STMR representation with LLMs significantly outperforms the direct input of RGB images with VLMs.
+TABLE VI: Ablation study on the different visual prompts for VLMs.
+VLM Input RGB STMR NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ SDTW/$\uparrow$
+GPT-4V $\checkmark$ 350.0 1.5 9.7 0.4
+$\checkmark$ 193.1 5.6 17.4 0.9
+GPT-4o $\checkmark$ 412.0 1.1 10.0 0.2
+$\checkmark$ 119.5 10.8 23.0 2.6
+TABLE VII: Ablation study on the grid size $r$ of STMR.
+Size NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ SDTW/$\uparrow$
+$10m$ 304.9 2.0 9.5 0.6
+$5m$ 119.5 10.8 23.0 1.9
+TABLE VIII: Ablation study on the state update strategies.
+Module NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ SDTW/$\uparrow$
+w/o state 289.7 9.0 22.1 1.7
+w state 119.5 10.8 23.0 1.9
 V CONCLUSIONS AND FUTURE WORK
-This paper addresses the challenging aerial VLN task by introducing an LLM-based end-to-end framework.
+This paper addresses the challenging aerial VLN task by introducing an LLM-based training-free framework.
 To enhance the spatial reasoning ability of LLMs, we propose Semantic-Topo-Metric Representation (STMR).
 STMR first integrates instruction-related landmarks and their locations into a top-down map, and subsequently transforms this map into a matrix representation containing semantic, topological, and distance metric information.
-Taking the proposed STMR as a part of prompts of the LLM, we significantly improve the UAV’s navigation capabilities.
-Our framework achieves state-of-the-art results on the AerialVLN-S dataset, demonstrating its effectiveness.
+Taking the proposed STMR as a part of the LLM prompts, we significantly improve the UAV’s navigation capabilities.
+Our framework achieves state-of-the-art results on the AerialVLN-S and Openfly dataset, demonstrating its effectiveness.
 In the future, we will explore flexible framework that will interact with humans to eliminate ambiguities in the instructions.

### v3

@@
 Aerial Vision-and-Language Navigation (VLN) is a novel task enabling Unmanned Aerial Vehicles (UAVs) to navigate in outdoor environments through natural language instructions and visual cues.
-However, it remains challenging due to the complex spatial relationships in outdoor aerial scenes.
+However, it remains challenging due to the complex spatial relationships in aerial scenes.
 In this paper, we propose a training-free, zero-shot framework for aerial VLN tasks, where the large language model (LLM) is leveraged as the agent for action prediction.
 Specifically, we develop a novel Semantic-Topo-Metric Representation (STMR) to enhance the spatial reasoning capabilities of LLMs.
-This is achieved by extracting and projecting instruction-related semantic masks of landmarks onto a top-down map, which contains location information about surrounding landmarks.
-The map is then transformed into a matrix representation with distance metrics, serving as the text prompt to the LLM for action prediction in response to the given instruction.
-Experiments conducted in real and simulation environments have successfully proved the effectiveness and robustness of our method, achieving at most 10.8% improvements (absolute) in Oracle Success Rate (OSR) on AerialVLN-S and Openfly dataset.
-I Introduction
-The Aerial Vision-and-Language Navigation (Aerial VLN) [liu_2023_AerialVLN] emerges as a groundbreaking task.
-It enables unmanned aerial vehicles (UAVs) to interpret natural language instructions and corresponding visual information to navigate in outdoor environments.
-This technology can eliminate the necessity for manual UAV operation by human pilots, clearly mitigating the barriers to human-UAV interaction and potentially benefitting search, rescue, and delivery tasks.
+This is achieved by extracting and projecting instruction-related semantic masks onto a top-down map, which presents spatial and topological information about surrounding landmarks and grows during the navigation process.
+At each step, a local map centered at the UAV is extracted from the growing top-down map, and transformed into a matrix representation with distance metrics, serving as the text prompt to LLM for action prediction in response to the given instruction.
+Experiments conducted in real and simulation environments have proved the effectiveness and robustness of our method, achieving absolute success rate improvements of 26.8% and 5.8% over current state-of-the-art methods on simple and complex navigation tasks, respectively.
+The dataset and code will be released soon.
+Introduction
+The Aerial Vision-and-Language Navigation (Aerial VLN) (Liu et al.
+2023c) emerges as a groundbreaking task.
+It enables unmanned aerial vehicles (UAVs) to interpret natural language instructions and visual information to navigate in outdoor environments.
+This technology can eliminate the necessity for manual UAV operation by human pilots, clearly mitigating the barriers to human-UAV interaction and potentially benefitting rescue, search, and delivery tasks.
+Figure 1: The pipeline to obtain STMR.
+(a) The observed RGB image, the corresponding segmented image, and the depth image.
+(b) Segmented images are projected into the top-down map gradually during the UAV flight, which captures the semantic and topological information of the environment.
+(c) The top-down map is further transformed into a 20x20 matrix representation with distance metrics for LLM reasoning.
 Recently, VLN tasks have been well-developed.
-Considering the powerful reasoning capability of large language models (LLMs) [ahn2022can, song2023llm, driess2023palm, huang2022inner], several VLN methods have started to use LLMs as agents to parse instructions and predict actions [zhou2024navgpt, shah2023lm, li2024tina, lin2024navcot].
-Specifically, existing methods attempt to describe the visual observations with text to enhance the LLM’s scene understanding ability.
-NavGPT [zhou2024navgpt] converted visual scene semantics into text description for the LLM to conduct high-level planning.
-MapGPT [chen2024mapgpt] introduced a topological graph described in language as text prompts to an LLM, guiding low-level planners to explore new nodes of a predefined graph of the environment.
-Figure 1: The pipeline to obtain STMR.
-Observation image, semantic mask, and depth image (a) can be aggregated in the top-down map (red dashed box) (b), which captures the Semantic-Topo-Metric information of the neighborhood of the current position, further generating a 20x20 matrix representation (c) for LLM reasoning.
-Although existing VLN methods have made significant progress, most of them are designed for indoor or ground-based outdoor environments and pay less attention to aerial VLN tasks.
-Practically, the rich semantics, complex spatial relationships, and the large-scale nature of outdoor 3D spaces make it challenging for existing VLN methods to adapt to aerial scenarios.
+Considering the powerful reasoning capability of large language models (LLMs) and vision language models (VLMs) (Ahn et al.
+2022; Song et al.
+2023; Driess et al.
+2023; Liu et al.
+2023a, 2024), several VLN methods have started to use LLMs or VLMs as agents to parse instructions and predict actions (Zhou, Hong, and Wu 2024; Shah et al.
+2023; Li, Chen, and Lin 2024; Lin et al.
+2024).
+Specifically, LLM-based methods attempt to describe the visual observations with text to enhance the LLM’s scene understanding ability (Zhou, Hong, and Wu 2024; Chen et al.
+2024b).
+VLM-based methods integrate the visual and text data and generates context-aware decisions, including (Zhang et al.
+2024c, b; Lee et al.
+2024; Sharshar et al.
+2025).
+Although existing LLM-based and VLM-based methods have made significant progress in indoor or ground-based outdoor environments, they still struggle to effectively encode the large-scale spatial information from the aerial view.
 As shown in  Fig.
-1 (a), the aerial scene can be quite complex, which may lead to incorrect reasoning for several reasons, such as an overemphasis on instruction-irrelevant objects, or a failure to capture the contextual relationships between different areas.
-Furthermore, the vast scale of aerial landscapes can further complicate navigation, making local-perspective-based VLN methods unable to comprehend the environment.
-Thus, the development of VLN methods that integrate both effective semantic information and precise spatial representations is urgently needed to enhance the adaptability and accuracy in aerial scenarios.
-To overcome these challenges, we propose a zero-shot LLM-based aerial VLN framework, which takes natural language instructions, RGB images, and depth images as input, and directly makes action predictions (e.g., forward 10 meters) through single-step reasoning updates.
+1, the aerial scene can be highly complex, which may lead to the overemphasis of instruction-irrelevant objects or the failure to capture the contextual relationships among different areas.
+Worse still, existing VLMs suffer from limitations in spatial reasoning when taking raw observations as input (Chen et al.
+2024a; Stogiannidis, McDonagh, and Tsaftaris 2025).
+Thus, there is an urgent need for the development of VLN methods that effectively integrate both semantic information and precise spatial representations to improve adaptability in aerial scenarios.
+To overcome these challenges, we propose a zero-shot LLM-based aerial VLN framework that encodes natural language instructions, RGB images and depth images as text prompt, and generates action predictions (e.g., go forward 10 meters) through LLMs directly.
 Specifically, a Semantic-Topo-Metric Representation (STMR) is designed for LLMs spatial reasoning.
-Leveraging the developments in open-vocabulary perception models [zhang2018review, jiang2018deep, minderer2022simple, cheng2024yolo], we first extract instruction-related landmarks and obtain corresponding semantic masks by perception models, i.e., Grounding Dino [liu2023grounding] and Tokenize Anything [pan2023tap].
-After that, semantic masks obtained from each step are gradually projected into a top-down map using depth and pose transformation as shown in Fig.
+Firstly, the instruction-related landmarks are extracted and corresponding semantic masks are generated via visual perception models, i.e., Grounding DINO (Liu et al.
+2023b) and Tokenize Anything (Pan et al.
+2023).
+Then, the semantic mask is projected into a top-down map as shown in Fig.
 1 (b).
-This projected map encompasses both the UAV’s travel trajectory and spatial information, i.e., locations of surrounding objects.
-To integrate rich visual information and topology into the text prompt while maintaining its simplicity, we separate the top-down map centered at the UAV’s current position into grids and substitute each grid with a semantic number.
-This process transforms the top-down map into a matrix representation, tailored to serve as a spatial prompt for the LLM.
+This top-down map encompasses both the UAV’s travel trajectory and spatial information, and it grows gradually during the navigation process.
+To encode the visual information into an LLM-friendly form, we firstly select a fixed-size region centered on the UAV in the top-down map as a local map, then separate the map into equally spaced grids and substitute each grid with a semantic number.
 As can be seen from Fig.
-1 (c), the matrix representation encompasses topological, semantic, and metric information, and is input to an LLM together with historical actions, sub-goal instructions, and other information to infer the next action.
-Compared with other learning-based and LLM-based methods, the proposed method effectively improves the accuracy of spatial reasoning.
+1 (c), the matrix representation encompasses topological, semantic, and metric information.
+It is input to an LLM together with historical actions and text instructions to infer the next action.
+Experiments have shown that the proposed method significantly enhances the capability of spatial reasoning, and achieves absolute success rate improvements of 26.8% and 5.8% over current state-of-the-art methods on simple and complex navigation tasks.
 Our contributions are summarized as follows:
   • To our knowledge, we design the first LLM-based training-free framework for the aerial VLN task, facilitating the development of UAV navigators.
-Without extra action planner, the proposed framework allows for easy integration.
+Without an extra action planner, the proposed framework allows for easy integration.
   • We propose the Semantic-Topo-Metric Representation (STMR), a unique matrix representation that encompasses topological, semantic, and metric information.
@@
   • Extensive experiments on the aerial VLN task demonstrate that the proposed method outperforms previous state-of-the-art methods by a large margin, establishing a strong baseline for future zero-shot aerial VLN tasks.
-II Related Work
-Vision-Language Navigation (VLN)
+Related Work
+Vision-Language Navigation (VLN).
 VLN aims to enable autonomous agents to navigate complex environments by understanding and executing natural language instructions based on visual context.
-Early VLN methods use sequence-to-sequence LSTMs to predict low-level actions [anderson2018vision] or high-level actions from panoramas [fried2018speaker].
-Several attention processes have been proposed [qi2020object, hong2020language, an2021neighbor] to enhance the process of learning visual textual correspondence.
-Reinforcement learning is also explored to improve policy learning [wang2018look, tan2019learning, wang2020soft].
-Besides, transformer-based architecture have shown superior performance to long-distance contextual information [hao2020towards, majumdar2020improving].
-More recent works [zhou2024navgpt, chen2024mapgpt] leverage the reasoning and dialogue capabilities of LLMs, achieving great progress.
-However, most of them are developed under a ground-based, discrete VLN setting, which simplifies navigation by traversing on a predefined graph on the ground.
-This limits the UAVs’ free movement space in the real world.
-Aerial VLN remains challenging due to the large-scale and complex environments.
-UAV navigation
-Unmanned Aerial Vehicle (UAV) navigation has seen a surge of interest over the past few decades.
-Guided with instructions, Blukis et al.
-[blukis2018following] use imitation learning for low-level velocity command prediction.
-[blukis2018mapping] predicts interpretable position-visitation distributions to guide UAV’s control actions, using a combination of supervised and imitation learning for efficient training.
-Misra et al.
-[misra2018mapping] decompose instruction execution in two stages, separately using supervised learning for goal prediction and policy gradient for action generation.
-AerialVLN [liu_2023_AerialVLN] contributes a much more challenging aerial VLN dataset focusing on outdoor aerial environments, and provides a look-ahead guidance method as the baseline.
-OpenFly [openfly] constructs a large-scale dataset across several high-quality scenes, and proposes a keyframe-aware VLN model.
+Early VLN methods use sequence-to-sequence LSTMs to predict low-level actions (Anderson et al.
+2018) or high-level actions from panoramas (Fried et al.
+2018).
+Several attention processes have been proposed (Qi et al.
+2020; Hong et al.
+2020; An et al.
+2021) to enhance the process of learning visual textual correspondence.
+Reinforcement learning is also explored to improve policy learning (Wang et al.
+2018; Tan, Yu, and Bansal 2019; Wang, Wu, and Shen 2020).
+Besides, transformer-based architecture have shown superior performance to long-distance contextual information (Hao et al.
+2020; Majumdar et al.
+2020).
+More recent works (Zhou, Hong, and Wu 2024; Chen et al.
+2024b) leverage the reasoning and dialogue capabilities of LLMs, achieving great progress.
+However, most of them operate in limited spaces on the ground.
+In contrast, aerial VLN remains challenging due to the large-scale and complex environments.
+UAV Navigation.
+Unmanned Aerial Vehicle (UAV) navigation has seen a surge of interest over the past years.
+Many earlier works (Blukis et al.
+2018a, b) using a combination of supervised and imitation learning for efficient training and low-level velocity prediction to guide UAV’s control actions.
+LINGUNET (Misra et al.
+2018) decomposes instruction execution into two stages, separately using supervised learning for goal prediction and policy gradient for action generation.
+AerialVLN (Liu et al.
+2023c) contributes a much more challenging aerial VLN dataset focusing on outdoor aerial environments, and provides a look-ahead guidance method as the baseline.
+Recent research (Wang et al.
+2024; Sautenkov et al.
+2025; Tian et al.
+2025) exploiting the powerful visual reasoning capabilities of VLM to address the challenges posed by UAV navigation scenes.
 Despite the progress, the generalizability and performance of these methods still require improvement.
-LLMs for robot planning and interaction
+LLMs for Robot Planning and Interaction.
 Most recently, LLMs have demonstrated impressive capabilities in understanding and reasoning.
 To leverage these capabilities, several promising methods have been proposed for applying LLMs in robotic systems.
-A few methods involve using LLM-generated rewards optimized in simulation to improve control [huang2023voxposer, yu2023language].
-Others utilize LLM-selected subgoals as an abstraction to enhance policies for navigation [dorbala2022clip, chen2023open] and manipulation [cui2022can, shridhar2022cliport].
-Additionally, research has explored the use of LLMs to generate executable code for control and perception primitives [singh2023progprompt, wu2023tidybot, liang2023code].
-Despite their potential, LLMs are still prone to confidently hallucinating outputs, such as referring to objects not observed in the scene [zeng2022socratic].
+A few methods involve using LLM-generated rewards optimized in simulation to improve control (Huang et al.
+2023; Yu et al.
+2023).
+Others utilize LLM-selected subgoals as an abstraction to enhance policies for navigation (Dorbala et al.
+2022; Chen et al.
+2023) and manipulation (Cui et al.
+2022; Li et al.
+2024).
+Additionally, research has explored the use of LLMs to generate executable code for control and perception primitives (Singh et al.
+2023; Liang et al.
+2023; Li et al.
+2025).
+Despite their potential, LLMs are still prone to confidently hallucinating outputs, such as referring to objects not observed in the scene (Zeng et al.
+2022).
 In order to alleviate the hallucinating phenomenon, we propose a matrix-based representation containing topological, semantic, and metric information for better prompting the LLMs.
-III Method
-In this paper, we introduce a novel zero-shot framework that leverages Large Language Models (LLMs) for action prediction in aerial VLN tasks.
+Method
+In this paper, we propose a zero-shot framework that leverages Large Language Models (LLMs) for action prediction in aerial VLN tasks.
 As shown in Fig.
-2, our framework mainly consists of three modules.
-The sub-goal extracting module decomposes language instructions into several sub-goals, facilitating step-by-step reasoning and navigation.
-The Semantic-Topo-Metric Representation (STMR) module represents the outdoor environment as a matrix containing semantic, topological, and metric information for the prompt to an LLM.
-Finally, the LLM planner outputs its current thoughts and actions by using the prompts from the aforementioned two modules along with the task description and history as inputs.
-III-A Problem Formulation
-In an aerial VLN task, the problem is formulated as a free-form language instruction guided navigation.
-At the beginning of each episode, the UAV is placed in an initial pose $P=[x,y,z,\phi,\theta,\psi]$, where $(x,y,z)$ denotes the UAV’s position and $(\phi,\theta,\psi)$ represents pitch, roll, and yaw of the UAV’s orientation.
+2, our framework consists of three modules.
+The sub-goal extraction module decomposes language instructions into several sub-goals, facilitating step-by-step reasoning and navigation.
+The Semantic-Topo-Metric Representation (STMR) module represents the outdoor environment as a matrix containing semantic, topological, and metric information, enhancing the spatial reasoning ability of LLMs.
+Finally, an LLM planner is designed, which takes sub-goal instructions, STMR, together with historical information and task description as input, and outputs its thoughts and predicted actions.
+Problem Formulation
+The aerial VLN task is formulated as a free-form language instruction guided navigation.
+At the beginning of each episode, the initial pose of a UAV is denoted as $P=[x,y,z,\phi,\theta,\psi]$, where $(x,y,z)$ is the UAV’s position and $(\phi,\theta,\psi)$ represents pitch, roll, and yaw of the UAV’s orientation.
 A natural language instruction $L$ is provided to specify the path that a UAV should follow.
-To achieve the navigation goal, the UAV considers both the instruction and visual perceptions, and predicts an action from the action space (right, left, up, down, forward, backward) with corresponding parameters at each time step $t$.
+To achieve the navigation goal, the UAV considers both the instruction and visual perceptions, and predicts an action from the action space (i.e., right, left, up, down, forward, backward) with corresponding value (e.g., 1-10m) at each time step $t$.
 Navigation ends when the UAV predicts a ‘Stop’ action or reaches a pre-defined maximum action number.
-Following [liu_2023_AerialVLN], the navigation is recognized as a success if the UAV stops at a location that is less than 20 meters from the target location.
-Figure 2: Our method consists of three modules, i.e.
-sub-goal extracting, Semantic-Topo-Metric Representation, and LLM planner.
-First, we prompt an LLM to break down long instructions and extract specific landmarks, which will be classified by perception models, generating corresponding semantic masks.
-Then we project these masks into a top-down map using the UAV-view pose and depth information.
-The map is further compressed into a matrix representation.
-Finally, we prompt LLM with well-designed instructions to enable effective reasoning.
-III-B Semantic-Topo-Metric Representation (STMR)
-Previous LLM-based VLN methods use natural language to describe current observations, or nodes and edges of a topological graph to model the spatial information of the environment.
-However, in open scenarios, simple directional words such as “next to” or “aside” are not enough to describe complex spatial relationships, which can easily introduce ambiguity into LLMs.
+Following (Liu et al.
+2023c), the navigation is considered successful if the UAV halts within 20 meters of the target location.
+Figure 2: Our method consists of three modules, i.e., Sub-Goal Extraction, Semantic-Topo-Metric Representation, and LLM planner.
+They are utilized to generate sub-goal instructions, spatial information representations, and UAV navigation actions, respectively.
+Semantic-Topo-Metric Representation (STMR)
+Previous LLM-based VLN methods use natural language to describe current observations, or a topological graph to model the spatial information of the environment.
+However, in open scenarios, simple directional words such as “next to” or “aside” are insufficient for describing complex spatial relationships, often leading to ambiguity in LLMs.
 To address this challenge, we introduce the STMR to enhance the spatial-aware reasoning capability of LLMs.
@@
 The details of STMR are presented as follows.
-III-B1 2D Visual Perception
+2D Visual Perception.
 Impressed by the powerful open-vocabulary detection capabilities of Grounding DINO $\operatorname{GD}(\cdot)$, as well as the captioning and segmentation capabilities of Tokenize Anything model $\operatorname{TA}(\cdot)$, we integrate these two models as our 2D visual preceptor, as illustrated in Fig.
@@
 Figure 3: 2D Visual Perceptor for the UAV.
-$\{m^{(t)},\ h^{(t)}\}=\operatorname{TA}\left({I_{t}^{R}},\ \operatorname{GD}\left({I_{t}^{R}},\ \operatorname{LE}({L})\right)\right).\$ (1)
-To improve the simplicity of the semantic masks and reduce the misleading of LLM reasoning caused by numerous open-vocabulary categories, we propose a text-matching method to mitigate over-classification.
+$\{m^{(t)},\ h^{(t)}\}=\operatorname{TA}\left({I_{t}^{R}},\ \operatorname{GD}\left({I_{t}^{R}},\ \operatorname{LE}({L})\right)\right).\\$ (1)
+To improve the robustness of the semantic masks and reduce the misleading of LLM reasoning caused by numerous open-vocabulary categories, we propose a text-matching method to mitigate over-classification.
 As shown in Fig.
 3, our method involves vectorizing the landmarks extracted from the instruction and the captions $h^{(t)}$ generated in each $I_{t}^{R}$.
-Then, we calculate the cosine similarity between these vectorized landmarks and captions using TF-IDF [qaiser2018text].
+Then, we calculate the cosine similarity between these vectorized landmarks and captions using TF-IDF (Qaiser and Ali 2018).
 If the similarity score exceeds the threshold $\tau$ (0.8), the landmark is classified as visible in the current view, and other irrelevant masks will be ignored.
 By implementing this strategy, we effectively simplify semantic masks and ensure the LLM’s reasoning focuses on relevant categories.
-III-B2 Sub-goal-driven top-down map
-Considering that the top-down view can better express spatial relationships than the first-person view, we further utilize the depth image $I_{t}^{D}$ to map semantic masks to 3D space, and then convert them into a top-down map.
-Specifically, by applying the above 2D visual perception process to the RGB image $I_{t}^{R}$, we obtain a segmented image with clearly identified objects and regions.
-Subsequently, the depth image $I_{t}^{D}$ is converted into a 3D point cloud, where each pixel is mapped to a corresponding 3D point $(X,Y,Z)$ based on its depth value and camera parameters:
+Sub-goal-driven Top-down Map.
+Considering that a top-down view better expresses spatial relations, we use the depth image to map semantic masks to the 3D space, and then project them into a top-down map.
+Note that the top-down map grows and updates during the UAV flight to support the navigation process.
+Specifically, by applying the above 2D visual perception process to the continuously observed RGB images, we obtain segmented images with identified objects and regions.
+Subsequently, the corresponding depth images are projected to a gradually growing 3D point cloud, where each pixel is mapped to a 3D point $(X,Y,Z)$ based on its depth value and camera parameters:
 $X=\frac{(u-c_{x})I_{t}^{D}(u,v)}{f_{x}},Y=\frac{(v-c_{y})I_{t}^{D}(u,v)}{f_{y}},$ (2)
@@
 where $(u,v)$ are the pixel coordinates, $(c_{x},c_{y})$ are the camera’s principal point coordinates, and $(f_{x},f_{y})$ are the focal lengths.
-The semantic labels from the segmented RGB image are mapped to the corresponding 3D point clouds, resulting in a point cloud with semantic information $(X,Y,Z,C_{i})$, where $C_{i}$ are the semantic categories.
-Then the 3D point cloud is partitioned into discrete voxels, where each voxel aggregates its point clouds as one semantic category using max pooling described in [chaplot2020learning].
-Considering that UAVs usually fly above ground objects, for a specific coordinate $(x,y)$, the semantic label of the top one in a column of voxels will be projected into the top-down map:
+The semantic labels from the segmented image are mapped to the corresponding 3D points, resulting in a point cloud with semantic information $(X,Y,Z,C_{i})$, where $C_{i}$ is the semantic category.
+Then the 3D point cloud is partitioned into discrete voxels, where each voxel aggregates its point clouds as one semantic category using max pooling.
+For a specific coordinate $(x,y)$, sometimes different categories of objects will appear at different heights.
+For example, there is ‘vegetation’ or ‘equipment’ on the roof of a ‘building’.
+Given that UAVs usually fly above the landmarks, the semantic label of the top one in a column of voxels will be projected into the top-down map:
 $\text{TopDownMap}(x,y)=\text{Voxel}(x,y,z_{top}),$ (4)
 where $z_{top}$ means the highest $z$ coordinate at location $(x,y)$ and $\text{Voxel}(x,y,z_{top})$ denotes the corresponding semantic label.
-So, we get a top-down map with semantic information.
+Consequently, we get a top-down map with semantic information.
 Notably, since the LLM prioritizes the landmarks in the current sub-goal, if a category contained in the sub-goal appears in the voxel at any $z$ coordinate, this category will be projected into the top-down map first.
@@
 $\text{TopDownMap}(x,y)=\begin{cases}C_{i},&C_{i}\text{ in sub-goal}\\ \text{Voxel}(x,y,z_{top}),&\text{otherwise}.\end{cases}$ (5)
-At each step, the position of a UAV will also be recorded as waypoints with special values and projected onto the top-down map.
-III-B3 Matrix representation
+Matrix Representation.
 We found that directly inputting the image-format top-down map to a vision-language model (VLM) often yields poor reasoning results.
-Alternatively, we process the visual map into a text-based matrix representation, which is input to an LLM to guide the navigation.
-Specifically, the UAV incrementally collects observations and gradually constructs the top-down map of the surrounding environment.
-Considering the perception quality of onboard sensors and the effective processing capacity of modern LLMs, we define a local map centered at the UAV’s current position as shown in Fig.
-1.
-Subsequently, we generate a matrix by partitioning the local map into a $20\times 20$ coordinate grid and applying semantic max pooling in each grid, making them finally contain one semantic category.
-Particularly, we grade all values that fall in each grid in the local map according to frequency, showing only the most frequent features as the final category of this grid.
-Each category such as building or car is discriminated by different number labels.
-Furthermore, the grid size $r$ (5m) is also input to the LLM as the metric information (see Table VII for ablation study on the grid size).
-In this case, the LLM can directly reason about positional relationships among each landmark.
-III-C Designing Text Prompt for LLM Planning
-Given that LLMs are well-trained in reasoning and planning, we complement their spatial perception and scene understanding capability with formatted prompt and action space.
-This prompt involves two components.
-First, the task description is prepared which includes the definition of the environment and the UAV, as well as the format of input and output.
-We require an LLM to leverage Chain-of-Thought reasoning [wei2022chain] after capturing the observation, thinking step by step in an order of observation-thought-planning-history to predict the next action.
-Regarding the input instruction, sub-goals are extracted by the LLM, and the landmarks given in the instructions will be updated in the task definition as the scene changes.
-Second, to improve the quality of prompts, we implemented a cyclic refinement process.
-This involves designing a new set of instructions for an LLM, including elements such as ‘history action’ and ‘next planning’ corresponding to the STMR module.
-The task description also emphasizes updating the status of planning, including three states, i.e., todo, in process, and completed, in order to minimize the problem of hallucinations that is common in LLM.
-The STMR is gradually updated in this iterative process as the UAV explores to improve the accuracy and relevance of the prompts.
-Example of LLM Prompt (simplified) [Task Description]You are an embodied unmanned aerial vehicle that navigates in the real world.
-You need to explore between some places marked and ultimately find the destination to stop.
-[Input Format] ‘Instruction’ is a global, step-by-step detailed guidance.
-‘History’ is your previously executed actions.
-‘Map’ is a top-down coordinate map, using a 20*20 grid M to represent coordinates, heading towards upper of the map.
-The value of the grid represent different landmarks.
-You are at [9,9] on the ’Map’ (marked as P).
-The distance between adjacent numbers is 5 meters in the real world.
-The numbers in the matrix M represent:
-[0:Unexplored 1:your past trajectory 2:road 3:building … ]. ‘Plan’ records previous long-term multi-step planning info that you can refer to now.
-[Output Format] ‘Thought’: your thoughts about the task, which may include your comprehension, surroundings, history, and etc.
-‘Observation’: the description of the map.
-‘Plan’: your updated plan.
-‘Action’: your next action.”
-Think step by step.
-First, describe your orientation and the landmark in your left, right and front.
-Second, according the ’map’, update your multi-step ‘Plan’.
-Each plan needs to follow a state word (TODO, In Process, Completed).
-Finally, give the specific “Action” according to the action format: Action: (right, left, lift, down, straight, back), (degree (0-90 degrees)), (distance (between 0-10 meters)).
-  A Simplified Input in the Cyclic Process
-Instruction
-: lift off and head straight across the water to the road…
-History: [lift off 5 meters…]
-Map: STMR
-Plan:
-1. Lift off 5 meters. (Completed)
-2. Turn 180 degrees. (Completed)
-3. Head straight across the water to the road. (Completed)
-4. Turn left to align the road. (In Process)
-5. (TODO)…
-IV Experiment
-TABLE I: Comparison on the validation seen of the AerialVLN-S. * denotes we train these methods by running their source codes.
-Validation Seen
-model NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ SDTW/$\uparrow$
+Alternatively, we process the visual map into an LLM-friendly text-based matrix representation, which effectively guides the navigation.
+In order to make LLMs aware of metric information, we define a 100m $\times$ 100m local map centered on the UAV’s current position from the entire top-down map, and divide it into $20\times 20$ coordinate grids with each grid covering a 5m $\times$ 5m area.
+The grid size is selected according to the size of the most common small landmarks (i.e., cars) to ensure them occupy a single grid, while larger landmarks can be represented by multiple grids.
+Subsequently, a text-based matrix is generated by applying semantic max pooling in each grid, where the most frequent category is selected as the semantic label.
+Each category, such as building or car, is identified by different number labels.
+Furthermore, the grid size $r$ (5m) is also input to the LLM as the metric information.
+Notably, the local map slides dynamically during the UAV flight, enhancing the LLM to understand the surrounding environment and reason about positional relationships among landmarks.
+LLM-based Navigator
+To improve the robustness of the LLM-based navigator, the prompt mainly consists of two components, i.e., task definition and STMR-based planner.
+Firstly, the task definition includes a description of the navigation task, as well as the input and output format.
+The input format specifies the size of the matrix representation and defines the mapping relationship between the semantic labels and matrix elements.
+The output format outlines the action space, i.e., (right, left, lift, down, straight, back), along with the maximum moving distance and turning range.
+Secondly, the STMR-based planner includes a text-based matrix representing contextual observation, sub-goal-based next-step planning, and history actions.
+The LLMs are required to leverage Chain-of-Thought (Wei et al.
+2022) reasoning after capturing the observation, thinking step-by-step in the order of observation-thought-planning-history to predict the next actions.
+During the navigation process, the status of each sub-goal is updated, consisting of three states, i.e., todo, in process, and completed, to further alleviate the issue of LLM hallucinations.
+The text-based matrix is progressively updated as the UAV navigates through the environment.
+The prompt examples can be found in the supplementary material.
+Experiment
+Dataset and Implementation Details
+Dataset.
+We conducted experiments using two test sets, i.e., the simple navigation test set collected by us and the complex navigation test set from the AerialVLN-S benchmark (Liu et al.
+2023c).
+Both test sets are derived from Microsoft AirSim plugins and Unreal Engine 4 scenes.
+Our simple test set consists of 1,000 trajectories that cover a range of common navigation scenarios, e.g., public transportation and urban environments.
+Each trajectory spans up to 100 meters and contains 1 or 2 landmarks as navigation references.
+This test set will also be released with the code.
+AerialVLN-S replicates real-world urban environments with over 870 object categories across diverse scenarios, including downtown cities, factories, parks, and villages.
+The flight trajectories are collected by AOPA-certified UAV pilots and curated by experts to ensure realism and navigational fidelity.
+For the AerialVLN-S benchmark, the trajectory is much longer (326.9 meters on average) with more landmarks and more complex instructions (83 words per instruction on average), causing an extremely challenging task.
+Besides, to verify the robustness and generalizability of the proposed method, we conduct real-world experiments in 10 outdoor scenes.
+Evaluation Metrics.
+To validate the effectiveness of our method, we utilize a comprehensive set of evaluation metrics following (Liu et al.
+2023c).
+We focus on several key aspects, i.e., Navigation Error (NE), quantifying the distance between the UAV’s stopping point and the actual destination; Success Rate (SR), measuring the proportion of navigations that successfully reach the destination within a 20-meter threshold; Oracle Success Rate (OSR), an idealized measure considering any point on the predicted trajectory that comes within 20 meters of the destination as a success; Normalised Dynamic Time Warping (SDTW), taking into account both the navigation success rate and the similarity between the predicted trajectory and ground truth.
+Implementation Details.
+Our framework is implemented in both a simulator and a real environment.
+In the simulation environments of AirSim and UE4, we evaluated the methods using a PC equipped with an Intel i9 12th-generation CPU and Nvidia RTX 4090 GPUs.
+For the real world environment, we test on a Q250 airframe, carrying an Intel RealSense D435i depth camera, and an NVIDIA Jetson Xavier NX running Ubuntu 18.04 as the onboard computer.
+This setup supports the construction of top-down maps and flight control.
+The deployment of large-scale models and communication with the onboard computer are managed by a mobile ground station equipped with high-performance computing capabilities (matching the simulation setup).
+For LLM reasoning, we utilize the online API of GPT-4o, employing the default parameters.
+Figure 4: Demonstration of the STMR in spatial reasoning.
+model NE/m $\downarrow$ SR/% $\uparrow$ OSR/% $\uparrow$ SDTW/% $\uparrow$
+Random 266.7 2.0 5.9 0.2
+Action Sampling 259.9 5.1 8.0 0.3
+Seq2Seq (Anderson et al. 2018) 130.1 25.5 57.1 10.6
+CMA (Hong et al. 2022) 113.6 29.8 73.2 10.5
+AerialVLN (Liu et al. 2023c) 116.4 37.7 71.7 12.7
+Navid (Zhang et al. 2024a) 99.5 44.1 79.9 16.6
+OpenUAV (Wang et al. 2024) 108.8 49.9 68.0 14.9
+Ours (training-free) 53.9 76.7 90.3 18.0
+Table 1: Overall performance comparisons on simple tasks.
+AerialVLN-S
+model NE/m $\downarrow$ SR/% $\uparrow$ OSR/% $\uparrow$ SDTW/% $\uparrow$
 Random 109.6 0.0 0.0 0.0
 Action Sampling 213.8 0.9 5.7 0.3
-LingUNet [misra2018mapping] 383.8 0.6 6.9 0.2
-Seq2Seq [anderson2018vision] 146.0 4.8 19.8 1.6
-CMA [CMA] 121.0 3.0 23.2 0.6
-AerialVLN [liu_2023_AerialVLN] 90.2 7.2 15.7 2.4
-Navid* [navid] 105.1 6.8 15.5 1.1
-Openfly* [openfly] 127.2 8.1 21.8 1.6
-Ours 96.3 12.6 31.6 2.7
-TABLE II: Comparison on the validation unseen of the AerialVLN-S. * denotes we train these methods by running their source codes.
-Validation Unseen
-model NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ SDTW/$\uparrow$
-Random 149.7 0.0 0.0 0.0
-Action Sampling 237.6 0.2 1.1 0.1
-LingUNet [misra2018mapping] 368.4 0.4 3.6 0.9
-Seq2Seq [anderson2018vision] 218.9 2.3 11.7 0.7
-CMA [CMA] 172.1 3.2 16.0 1.1
-AerialVLN [liu_2023_AerialVLN] 127.9 5.1 10.5 1.4
-Navid* [navid] 106.9 6.1 12.3 0.7
-Openfly* [openfly] 113.8 7.6 18.2 1.5
-Ours 119.5 10.8 23.0 1.9
-TABLE III: Comparison on OpenFly dataset unseen split.
-OpenFly Unseen
-model NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ SDTW/$\uparrow$
-Random 301m 0.1 0.1 0
-Seq2Seq[anderson2018vision] 291m 0.1 3.2 0.1
-CMA[CMA] 268m 0.7 6.6 0.1
-AerialVLN[liu_2023_AerialVLN] 227m 2.3 19.8 1.3
-Navid[navid] 200m 2.6 16.9 2.2
-OpenFly[openfly] 198m 5.1 27.3 2.0
-Ours 205m 8.0 31.0 2.4
-IV-A Dataset and Implementation Details
-IV-A1 Dataset
-We conduct experiments using two datasets to assess aerial VLN task performance, i.e., AerialVLN-S benchmark from [liu_2023_AerialVLN] and the recently introduced OpenFly dataset [openfly].
-The AerialVLN-S dataset simulates real-world urban environments with over 870 object categories across diverse scenarios such as downtown cities, factories, parks, and villages.
-It contains 8,446 expert-curated flying paths recorded by AOPA-certified UAV pilots, ensuring trajectory realism and navigational fidelity.
-For evaluation, we utilize 12 scenes from its validation set and 5 scenes from the validation unseen set, as the official test set remains inaccessible.
-To enable broader comparative analysis, we introduce the OpenFly unseen dataset, which provides 1,200 test trajectories uniformly sampled from more diverse scenarios.
-We compare with several state-of-the-art VLN models on both benchmarks, using OpenFly’s test unseen trajectories for generalization assessment.
-This dual evaluation strategy rigorously tests the performances of these works, establishing comprehensive baselines for aerial VLN research.
-IV-A2 Evaluation metrics
-To validate the effectiveness of our method, we utilize a comprehensive set of evaluation metrics following [liu_2023_AerialVLN].
-We focus on several key aspects, i.e., Navigation Error (NE) quantifying the distance between the UAV’s stopping point and the actual destination, Success Rate (SR) which measures the proportion of navigations that successfully reach the destination within a 20-meter threshold, Oracle Success Rate (OSR), an idealized measure considering any point on the trajectory that comes within 20 meters of the destination as a success, and Success rate weighted by Normalised Dynamic Time Warping (SDTW), which takes into account both the navigation success rate and the similarity between the real paths.
-IV-A3 Implementation details
-Our framework is implemented in both a simulator and a real environment.
-In the simulation environments of AirSim and Unreal Engine 4 (UE4), we evaluated the methods using a PC equipped with an Intel i9 12th-generation CPU and Nvidia RTX 4090 GPUs.
-For the real world environment, we test on a Q250 airframe, carrying an Intel RealSense D435i depth camera, and an NVIDIA Jetson Xavier NX running Ubuntu 18.04 as onboard computer.
-All methods are run on an external PC (same as the simulation setup) communicating with the onboard computer via a UDP network.
-For LLM reasoning we use the online API of GPT-4V and GPT-4o, where the default parameters are employed.
-IV-B Experimental Results
-IV-B1 Quantitative results in simulator
-Table I and  Table II compare our method with several other works on the AerialVLN-S dataset.
-Random and Action Sampling respectively mean sampling actions randomly and according to the action distribution of the training set.
-LingUNet [misra2018mapping] assumes an agent can see the destination from the beginning, which is transformed into a step-wise paradigm by [liu_2023_AerialVLN].
-Seq2Seq [anderson2018vision] is a learning-based recurrent policy, directly predicting the next action.
-CMA is a baseline model in [liu_2023_AerialVLN], considering attention between multi-modalities.
-LAG is the main method in [liu_2023_AerialVLN], designing a look-ahead guidance to adjust paths.
-Navid [navid] is a VLM-based method, achieving great success in the indoor VLN task.
-OpenFly [openfly] proposes a keyframe-aware aerial VLN model, taking key historical observations into account.
-In comparison, our method significantly outperforms others in most metrics, demonstrating the superiority of our method.
-Notably, our method is training-free, while others are trained using corresponding data.
-Table III shows the generalization results of different methods on the OpenFly Unseen dataset, where our method also achieves better performance than other works.
-We found that instead of processing the entire long instruction and visual observations at each time step, the formatted decomposed sub-goal is more conducive to the fine-grained alignment of each action and landmark.
-Moreover, UAV has a better understanding of its own spatial position and other semantic constraints under the guidance of our STMR, which has obvious advantages in long instruction environments and leads to more robust execution.
-Aerial VLN is a challenging task that has not been fully studied, and the overall performance needs to be improved.
-We believe that our work can bring some insights into this field.
-IV-B2 Quantitative results in real environment
-To test whether our solution works well in the real world, we collect 10 outdoor scenes including street scenes and forests with ground-truth lengths ranging from 50m to 500m.
-Then, we evaluate our method using a real UAV and call cloud-hosted LLMs to navigate in these challenging environments.
-Not only pretrained VLN models but also other LLM-based VLN methods, i.e., MapGPT [chen2024mapgpt] and NavGPT [zhou2024navgpt], are employed and compared in this experiment.
-Notably, MapGPT and NavGPT rely on predefined topological maps to predict the next navigation waypoint, we manually define candidate waypoints for them in this experiment.
-From demonstration snapshot Fig.
-6, we can see that our method can align visual and textual landmarks as well as understand commands.
-Finally, the UAV reaches the destination successfully.
-Table IV also shows that the proposed method achieves a much better performance than both pretrained VLN methods and training-free LLM-based methods.
-However, we can see that LLM-based methods including ours need more computation time.
+LingUNet (Misra et al. 2018) 383.8 0.6 6.9 0.2
+Seq2Seq (Anderson et al. 2018) 146.0 4.8 19.8 1.6
+CMA (Hong et al. 2022) 121.0 3.0 23.2 0.6
+AerialVLN (Liu et al. 2023c) 90.2 7.2 15.7 2.4
+Navid (Zhang et al. 2024a) 105.1 6.8 15.5 1.1
+OpenUAV (Wang et al. 2024) 102.8 6.3 17.6 2.0
+Ours (training-free) 96.3 12.6 31.6 2.7
+Table 2: Comparison on the validation seen set of AerialVLN-S.
+Experimental Results
+Baseline Models
+  • $\bullet$
+    Rule-based Methods.
+Rule-based methods include Random and Action Sampling.
+In the Random method, the agent randomly chooses actions at each step and continues until either the ‘stop’ action is chosen or the maximum number of steps is reached.
+The Action Sampling method, on the other hand, first analyzes the statistical properties of the dataset and then samples actions based on the action distribution.
+  • $\bullet$
+    Learning-based Methods.
+Learning-based methods include mainstream aerial navigation approaches like LingUNet, Seq2Seq and CMA (Liu et al.
+2023c), as well as the state-of-the-art method Navid (Zhang et al.
+2024a) and OpenUAV (Wang et al.
+2024).
+Different from other baselines, Navid is a method for the indoor VLN task, so we retrain it on the original AerialVLN dataset.
+  • $\bullet$
+    LLM-based Methods.
+For LLM-based methods, we compare two works for the indoor VLN task, i.e., MapGPT (Chen et al.
+2024b) and NavGPT (Zhou, Hong, and Wu 2024).
+To ensure fairness, all these methods are evaluated using the GPT-4o with the same setting.
+Quantitative Results in Simulator.
+We first analyze performance on the simple navigation tasks.
+The scenes mainly tests the agent’s ability to locate landmarks and predict actions correctly.
+Even a short instruction like “fly along the road and stop near the intersection” requires the UAV to accurately understand the scene, recognize the destination, and stop in a proper location.
+As shown in Table 1, our method outperforms other works by a large margin in terms of SR (+26.8%) and OSR (+22.3%).
+In addition, our method exhibits a significantly lower navigation error (NE) compared to the other methods, which is reduced by 45.6 meters on average compared with the second-place method, Navid.
+These indicate that the UAV has a better understanding of its own spatial position and semantic constraints under the guidance of our STMR.
+In Table 2, the proposed method consistently achieves superior performance on the AerialVLN-S dataset.
+As the task complexity increases significantly with much longer trajectories and instructions, the success rates of all methods experience a notable decline.
+The results in Table 2 indicate that Aerial VLN remains a challenging task having not been fully studied, and there is still room for improvement in overall performance.
+We believe that our work can bring valuable insights into this field.
+Quantitative Results in Real Environment.
+To test the performance of our solution in the real world, we collected 10 outdoor scenes, including street scenes and forests, with ground-truth lengths ranging from 50m to 200m.
+Then, We evaluate our method using a real UAV, leveraging cloud-hosted LLMs to navigate through these challenging environments.
+In this experiment, we employ and compare not only pretrained VLN models but also other LLM-based VLN methods, i.e., MapGPT (Chen et al.
+2024b) and NavGPT (Zhou, Hong, and Wu 2024).
+Notably, MapGPT and NavGPT rely on predefined topological maps to predict the next navigation waypoint.
+In this experiment, we manually define candidate waypoints for both methods.
+As shown in the demonstration snapshot in Fig.
+6, our method effectively aligns visual and textual landmarks, understands commands, and successfully guides the UAV to its destination.
+Table 3 also shows that the proposed method achieves a much better performance than both pretrained VLN methods and training-free LLM-based methods.
+However, we can see that LLM-based methods, including ours, need more computation time.
 This is mainly because of the latency of the LLM API.
 We believe that with the development of LLMs and hardware computing power, the latency will gradually decrease.
-Figure 4: Demonstration of the STMR in spatial reasoning.
 Figure 5: Failure cases of our method.
-TABLE IV: Quantitative results in real environment. The last column shows the average runtime of aerial VLN procedures.
-model SR/$\uparrow$ OSR/$\uparrow$ Time (s)
-AerialVLN [liu_2023_AerialVLN] 0 10 82
-Navid [navid] 10 10 75
-Openfly [openfly] 20 30 63
-MapGPT [chen2024mapgpt] 20 20 176
-NavGPT [zhou2024navgpt] 10 20 150
+model SR/% $\uparrow$ OSR/% $\uparrow$ Time (s)
+AerialVLN (Liu et al. 2023c) 0 10 82
+Navid (Zhang et al. 2024a) 10 10 75
+MapGPT (Chen et al. 2024b) 20 20 176
+NavGPT (Zhou, Hong, and Wu 2024) 10 20 150
 Ours 40 70 117
-Figure 6: Qualitative results in real environment. Short-range movements of the same action are merged into a single long-range movement, as shown in the image sequence.
-IV-B3 Case analysis
+Table 3: Quantitative results in real environment. The last column shows the average runtime of aerial VLN procedures.
+Figure 6: Visualization of successful and failed examples in the real environments.
+Short-range movements of the same action are merged into a single long-range movement, as shown in the image sequence.
+The second row illustrates a failed case, the UAV mistakenly executed the “go straight” action because the visual perceptor failed to recognize the road in the fifth image.
+Case Analysis.
 Fig.
 4 demonstrates the effectiveness of our STMR in helping spatial reasoning, where the left part shows an intermediate STMR during the navigation process, and the right part shows simplified prompt, questions, and reasoning answers.
-From Fig.
-4 we can see that an LLM can understand and describe the surrounding environment through the STMR.
+As shown in Fig.
+4, an LLM can understand and describe the surrounding environment through the STMR.
 It can even make accurate judgments about objects or areas at a distance.
-Such an ability is extremely helpful for aerial VLN in vast outdoor environments.
-We also summarize two most common failures caused by incorrect planning or execution.
-The first typical failure is the misunderstanding of ambiguous instructions.
-For example, as shown in the left part of Fig.
-5, sometimes there are a lot of continuous commands without any landmark, such as “turn left, then move right, then go straight”, which have no objects for an LLM to refer to.
-This case will directly lead to the repeated execution of an action.
-The second typical error is caused by the inaccuracy of the visual perception.
-Although modern perception models show strong capability, they still need to be improved in identifying objects from different views.
+This ability is particularly valuable for aerial VLN in vast outdoor environments.
+Fig.
+5 illustrates two of the most common failure cases caused by incorrect planning or execution.
+The first typical failure arises from the misunderstanding of ambiguous instructions.
+there are many continuous commands without any landmarks, such as turn left, then move right, then go straight, which lack landmarks for the LLM to reference.
+This ambiguity often leads to the repeated execution of an action.
+The second typical error is caused by the inaccuracy of visual perception.
+Although modern perception models show strong capability, they still require improvement in identifying objects from different views.
 As a result, the key landmark may not be mapped to the STMR, causing the proper action not to be completed as illustrated in the right part of Fig.
-5.
-This observation aligns with our real-world experiment.
-TABLE V: Ablation study on different spatial representations.
+5 and the second row of Fig.
+6, showcasing the challenges posed by complex environments.
 Method Validation Unseen
-NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ SDTW/$\uparrow$
+NE/m $\downarrow$ SR/% $\uparrow$ OSR/% $\uparrow$ SDTW/% $\uparrow$
 Topo 203.3 4.9 12.8 1.8
 Metric 165.0 6.1 13.7 2.2
-Ours 87.1 19.4 26.2 5.7
-IV-C Ablation Studies
-We carry out comprehensive ablation studies to assess the core designs of the proposed method.
-All ablation experiments are conducted on the randomly selected 100 samples from the AerialVLN-S dataset.
-Corresponding results are depicted in Table V,  Table VI,  Table VII, and Table VIII.
-IV-C1 Different types of spatial information
+Ours 88.7 15.0 28.0 3.6
+Table 4: Ablation study on different spatial representations.
+Ablation Studies
+We conduct comprehensive ablation studies to assess the core components of the proposed method.
+We randomly sample 100 samples from the unseen validation dataset of AerialVLN-S and perform all the ablation experiments.
+Corresponding results are depicted in Table 4,  Table 5,  Fig.
+7, and Table 6.
+Different Types of Spatial Information.
 To demonstrate the spatial representation ability of STMR, we further compare it with other prompting formats as an LLM’s spatial information prompt.
-In Topo format (row 1 of Table V), we maintain a linguistically formed map that captures the topological relationships between different nodes.
+In Topo format (row 1 of Table 4), we maintain a linguistically formed map that captures the topological relationships between different nodes.
 Each node records textual descriptions of visual observations, and the connectivity between nodes is described using textual prompts.
-For example, “Place 1 is connected with Places 2, 4, 0, 3”.
-In Metric format (row 2 of Table V), we arrange the visual observations from 8 different directions in a clockwise order relative to the agent’s current orientation and concatenate them into a single prompt.
-Each vision observation records the direction and distance of the landmark, for example, “a white building in the left front 10 meters away”.
-It is shown that row 2 improves SR by 1.2% (absolute) over row 1 but still has poorer navigation performance, since for some landmarks such as roads and rivers, UAVs still cannot obtain region-level spatial information based on a certain point distance description.
-Our STMR combining the semantic, topological, and metric information significantly enhances the UAV’s exploration capability and improves the accuracy in SR and OSR by over 10%.
-IV-C2 Direct visual input vs STMR
-Due to the powerful multimodal representation capabilities of GPT-4V and GPT-4o, they are capable of direct reasoning with visual and textual prompts.
+For example, Place 1 is connected with Places 2, 4, 0, 3.
+In the Metric format (row 2 of Table 4), we arrange the visual observations from 8 different directions in a clockwise order relative to the agent’s current orientation and concatenate them into a single prompt.
+Each visual observation records the direction and distance of the landmark, for example, a white building in the left front 10 meters away.
+It is shown that row 2 improves SR by 1.2% (absolute) over row 1 but still demonstrates poorer navigation performance.
+This is because, for certain landmarks like roads and rivers, UAVs still cannot obtain region-level spatial information based solely on point-distance descriptions.
+In contrast, our STMR combining the semantic, topological, and metric information, significantly enhances the UAV’s exploration capability and improves OSR accuracy by over 10%.
+Direct Visual Input vs STMR.
+Due to the powerful multimodal representation capabilities of GPT-4V and GPT-4o, they can directly reason with both visual and textual prompts.
 However, it is challenging for them to make correct VLN decisions from sequential images.
-This may be because of the inability of VLMs in spatial-aware reasoning.
-As shown in Table VI, the proposed STMR representation with LLMs significantly outperforms the direct input of RGB images into VLMs.
-IV-C3 Grid size of STMR
-[…]
-IV-C4 State update of sub-goal
-To ensure that each sub-goal can be executed despite environmental changes, in each iteration, the LLM does not change the original path plan, but updates the status of the current sub-goal including todo, in process, and completed.
-We compare this strategy with updating the entire multi-step path planning (row 1 of Table VIII) according to the observation in each iteration.
-This strategy does not notably enhance the Oracle Success Rate (OSR), showing that the proposed STMR already furnishes enough capacity for navigating the environment comprehensively.
-Nonetheless, it positively impacts the decision-making process, resulting in an increase in the success rate (SR) from 9.0% to 10.8%.
-TABLE VI: Ablation study on the different visual prompts for VLMs.
-VLM Input RGB STMR NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ SDTW/$\uparrow$
+This may be attributed to the inability of VLMs in spatial-aware reasoning.
+As shown in Table 5, the proposed STMR representation with LLMs significantly outperforms the direct input of RGB images into VLMs.
+It demonstrates that LLMs can get better spatial reasoning abilities when provided with proper visual encoding.
+Grid Size of STMR.
+In Fig.
+7, we demonstrate the necessity of maintaining a trade-off between grid resolution and the total number of grid cells in STMR representation.
+Set the resolution to 5m may not allow for as precise spatial inference as 2m, but it is enough for outdoor spatial reasoning.
+State Update of Sub-goal.
+To ensure that each sub-goal can be executed despite environmental changes, the LLM does not change the original path plan in each iteration.
+Instead, it updates the status of the current sub-goal, including todo, in process, and completed.
+We compare this strategy with one that updates the entire multi-step path planning (row 1 of Table 6) based on the observation in each iteration.
+This strategy does not notably enhance the OSR performance, indicating that the proposed STMR already furnishes enough capacity for navigating the environment comprehensively.
+Nonetheless, it does positively impact the decision-making process, leading to an increase in the success rate (SR) from 9.0% to 15.0%.
+VLM Input RGB STMR NE/m $\downarrow$ SR/% $\uparrow$ OSR/% $\uparrow$ SDTW/% $\uparrow$
 GPT-4V $\checkmark$ 350.0 1.5 9.7 0.4
-$\checkmark$ 193.1 5.6 17.4 0.9
+$\checkmark$ 112.5 10.2 22.4 2.1
 GPT-4o $\checkmark$ 412.0 1.1 10.0 0.2
-$\checkmark$ 119.5 10.8 23.0 2.6
-TABLE VII: Ablation study on the grid size $r$ of STMR.
-Size NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ SDTW/$\uparrow$
-$10m$ 304.9 2.0 9.5 0.6
-$5m$ 119.5 10.8 23.0 1.9
-TABLE VIII: Ablation study on the state update strategies.
-Module NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ SDTW/$\uparrow$
+$\checkmark$ 88.7 15.0 28.0 3.6
+Table 5: Ablation study on the different visual prompts for VLMs.
+Figure 7: Ablation study on different STMR grid sizes.
+Module NE/m $\downarrow$ SR/% $\uparrow$ OSR/% $\uparrow$ SDTW/% $\uparrow$
 w/o state 289.7 9.0 22.1 1.7
-w state 119.5 10.8 23.0 1.9
-V CONCLUSIONS AND FUTURE WORK
-This paper addresses the challenging aerial VLN task by introducing an LLM-based training-free framework.
-To enhance the spatial reasoning ability of LLMs, we propose Semantic-Topo-Metric Representation (STMR).
+w state 88.7 15.0 28.0 3.6
+Table 6: Ablation study on the state update strategies.
+Conclusions
+This paper addresses the challenging aerial VLN task by proposing an LLM-based training-free framework.
+To enhance the spatial reasoning ability of LLMs, we design Semantic-Topo-Metric Representation (STMR).
 STMR first integrates instruction-related landmarks and their locations into a top-down map, and subsequently transforms this map into a matrix representation containing semantic, topological, and distance metric information.
 Taking the proposed STMR as a part of the LLM prompts, we significantly improve the UAV’s navigation capabilities.
-Our framework achieves state-of-the-art results on the AerialVLN-S and Openfly dataset, demonstrating its effectiveness.
-In the future, we will explore flexible framework that will interact with humans to eliminate ambiguities in the instructions.
+Our framework achieves state-of-the-art results on both simple navigation tasks and complex navigation tasks, demonstrating its effectiveness and robustness.
+Limitations and Future Work
+While the proposed STMR significantly improves LLM’s spatial understanding ability, it still struggles to handle very precise distances in 3D space.
+Additionally, pretrained perception models sometimes introduce errors under different observation angles and distances, causing difficulties for subsequent navigation.
+Future avenues of improvement include using the primary viewpoint to assist spatial perception, or designing more robust perception strategies.